### v2

# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators.
Keywords— Endogeneity, Super Learner, Identification, Control Function, Panel data
1 Introduction
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders, and one some of them are time-invariant. In this setup, it is common to assume that the structural is linear or to exploit a linear projection in the reduced form equation. However, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to problems due to weak instruments since the strength of the linear relationship might not be high. This paper proposes to exploit the nonlinearties in the relationship between the endogenous covariate and the instrumental variables as a way to avoid problems of weak instruments due to nonlinearities.  [^1] Instruments are weak when they are weakly correlated to the endogenous covariate. In a cross-sectional setup, weak IVs causes that the median of the 2SLS estimator is biased towards the OLS estimator, 2SLS t-test presents size inflation (Keane and Neal (2024)). In panel data, Within-2SLS estimator is consistent when the time dimension grows (Cai et al. (2012)).
An example is the effect of air pollution exposure on children’s educational outcomes, where both variables are influenced by unobserved confounders that may be either time-varying or time-invariant. For instance, neighborhood characteristics and family income are often unobserved in publicly available data. While neighborhood characteristics typically remain constant over time, family income fluctuates. Both factors are relevant to a child’s academic achievement and exposure to air pollution, leading to endogeneity in pollution exposure. To address this, fixed effects can be used to control for unobserved neighborhood characteristics, while an instrumental variable can help account for time-varying unobservables. A commonly used instrumental variable is wind direction, which has a well-established nonlinear relationship with air pollution levels (Zabrocki et al., 2022).
This paper studies a triangular structural panel data model where the reduced form equation is nonlinear and individual specific effects are present. It proposes the use of machine learning methods for the estimation of the nonlinearities in the model. For simplicity consider the following triangular model without covariates for $i\in\{1,...,N\}$, $t\in\{1,...,T\}$:
$y_{it}=x_{1it}\beta_{1_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad with\quad E[\varepsilon_{it}|z_{i},\alpha_{i,y}]=0,$ (1)
$x_{1it}=g_{o}(z_{it})+\alpha_{i,1x}+u_{it},\quad with\quad E[u_{it}|z_{i},\alpha_{i,1x}]=0,$ (2)
$\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad with\quad E[\omega_{it}|u_{i},z_{i},\alpha_{i,y}]=0,$ (3)
[…]
The model comprises a structural equation (1) with an endogenous regressor $x_{1it}$ that presents a non-linear relationship with the available instrumental variable $z_{it}$ (2). Both equations present individual effects that represent time-invariant unobserved confounders. In this setup, conditioning on the individual effects allows us to control for time-invariant unobserved confounders. But, estimating all the individual effects leads to incidental parameter bias as the number of parameters to estimate grows with the sample size. Then, in order to eliminate the individual effects, we propose to either first-difference the model or perform a within transformation (obtain the deviations of the original variables from their individual mean). On the other hand, the presence of time-varying unobserved confounders is modeled through a linear relationship between the error terms in both equations (3). We deal with this issue by exploiting available exclusion restrictions (instrumental variables) along with a Control Function approach on the transformed model.
To estimate the causal parameter of interest $\beta_{1_{o}}$, we provide a two-step estimator to estimate the model with a linear structural equation. In the first step, we use a super learner (Van der Laan et al., 2007) to learn the nonlinear relationship of the first-differenced endogenous variable with the instrumental variables and the observable confounders. After this, we obtain the residuals of the first-differenced endogenous variable. In the second step, we use the estimated residuals in the first stage as a control function for the endogenous variable in the structural equation using two different score functions. We call this estimator Super Learner Control Function estimator.
Using a Monte Carlo experiment, we show that the estimators perform well provided that we can accurately learn the nuisance parameter in the first stage. The development of the statistical properties of the proposed estimator is ongoing, and further details are available upon request from the author.
Our paper contributes to the literature on the use of machine learning to estimate panel data models. We are the first to use super learning to estimate a panel data model with an endogenous covariate (due to unobserved confounders or measurement error) and additive individual fixed effects using a control function approach and exploiting nonlinearities. In addition, this paper is the first to use either First-Difference or Within transformation to take care of the individual specific effects while using machine learning methods. In contrast, these transformations have been used in the nonparametric panel data literature. The available papers using machine learning to estimate panel data models focus on exogenous covariates (time-varying confounders are observed) and the presence of individual fixed effects. Klosin and Vilgalys (2022) proposes a debiased LASSO estimator to estimate heterogeneous treatment effects of a continuous treatment, which is modeled using a structural nonparametric panel data model with additive fixed effects and strictly exogenous covariates. They account for the individual specific effects by taking the first-difference of the original model. Semenova et al. (2023) focuses on a structural equation model for panel data allowing for heterogenous treatment effects with unconfoundedness on observables (sequentially exogenous covariates) and individual specific fixed effects. In order to deal with the individual fixed effects, the authors use a Mundlak approach. They provide an algorithm composed of three steps. In the first step, using cross-fitting they estimate the nuisance functions of the structural and treatment equations to obtain the residuals. In the second step, they estimate the conditional average treatment effect using lasso estimators. In the third step, they provide debiased inference.
This paper also contributes to the literature on weak instruments in panel data by being the first one to consider the problem of weak instruments due to nonlinearities, and to use super learning with control function approach as a solution. The available papers on weak instrumental variables in panel data focus on the asymptotic properties of the Within 2SLS estimator showing that it is consistent even with weak instruments when the time dimension grows to infinity. Cai et al. (2012) show that the Within 2SLS is consistent when the individual and time dimension grow to infinity. They also present asymptotical pivotal tests (Anderson-Rubin test, Kleibergen test) for longitudinal data. Cai et al. (2012) provide conditions for consistency of the Within 2SLS estimator.
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5 we describe the identification strategy. In Section 6, we describe the control function approach. In Section 7, we present the simulation experiment and its results, in Section 8 we present the results of the empirical application, and in Section 9 we present the conclusions of the paper.
2 Other Related literature
This paper is related to the literature on semiparametric panel data models. Rodriguez-Poo and Soberon (2017) explains that one can either use a first-differencing approach or a profiling technique to estimate a semi-parametric model with an additive separable individual specific effect and a disturbance term. Li (1996) explains that kernel estimation of the conditional expectation of the first-differenced endogenous regressor is not feasible when the dimension of the conditioning set is larger than 5. As a solution, Baltagi et al. (2002) presents an additive estimation of the first-differenced non-linear function.
In addition, this paper is related to the literature on orthogonal/debiased estimation, which focuses on providing $\sqrt{N}$ consistent estimators of low-dimensional parameters in the presence of high-dimensional nuisance parameters. Chernozhukov et al. (2018) defines an orthogonal score as the one that presents a vanishing Gateaux derivative with respect to the nuisance parameters when evaluated at the true finite-dimensional parameter values. The orthogonal score is closely related to the score proposed by Robinson (1988).
Finally, our work relates to the literature of mixed effects models and machine learning. Our model is related to a mixed effects model assuming that covariates are exogenous conditioning on the cluster random effects. Emmenegger and Bühlmann (2023) uses double machine learning to estimate a partially linear mixed-effects model. Young and Shah (2024) proposes a sandwich boosting for accurately estimating partially linear models for grouped data.
3 Motivating examples
As mentioned before, measuring the causal effect of air pollution on health and educational outcomes is not straightforward as it is not randomly assigned. More specifically, time-varying unobservables such as family income might affect air pollution and health or educational outcomes. In addition, unobserved time-invariant variables, such as neighborhood characteristics or genetics, might also be present. While time-invariant unobservables could be dealt with by including individual fixed effects in the model, the presence of time-varying unobservable confounders requires instrumental variables. In this setup, wind direction is a commonly used instrumental variable. While it is well documented that the relationship between air pollution and wind direction is nonlinear (Zabrocki et al., 2022), it is common to see that the nonlinearity is ignored or modeled with a parametric assumption.
3.1 Air pollution and educational outcomes
Gilraine and Zheng (2022) studies the effect of air pollution on educational outcomes. They specify the following linear regression for subject $s$, cohort $c$, district $d$, year $t$:
$y_{sdct}=\beta PM2.5_{dt}+X_{sdct}^{\prime}\gamma+W_{dt}^{\prime}\eta+\omega_{s}+\alpha_{d}+\theta_{c}+\lambda_{t}+\epsilon_{scdt},$ (4)
where $y_{sdct}$ represents the average grade of subject $s$, cohort $c$, district $d$ at year $t$. $PM2.5_{dt}$ represents the average level of $PM2.5$ at district $d$ in year $t$, $X_{sdct}$ is a vector of control variables, $W_{dt}$ is a vector of weather variables, $\omega_{s}$ represents subject specific effects, $\alpha_{d}$ represents district fixed effects, $\theta_{c}$ represents cohort fixed effects, $\lambda_{t}$ are time fixed effects. and $\epsilon_{scdt}$ is the disturbance term.
As mentioned before, $PM2.5_{dt}$ is not randomly assigned and subject to measurement error. Thus, the authors use year-to-year changes in coal production and a shift-share instrument equal to the interaction of fuel shares used for nearby production with national growth rates along with a linear first-stage regression.
As an alternative, following Deryugina et al. (2019) we could use wind direction as an instrument of air pollution along with a nonlinear first-stage equation as follows:
$PM25_{dt}=h(WD_{dt},X_{dt}^{k})+\gamma_{d}+\varepsilon_{dt},$ (5)
with $WD_{dt}$ is the average wind direction in district $d$ in year $t$, and $\gamma_{d}$ district specific effects.
3.2 Air pollution and health outcomes
Deryugina and Reif (2023) studies the effect of air pollution on mortality rates. More precisely, they specify the following linear regression for county $c$, day $d$, month $m$ and year $y$:
$y_{cdmy}^{k}=\beta PM2.5_{cdmy}+X_{cdmy}^{\prime}\gamma+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\epsilon_{cdmy},$ (6)
where $y_{cdmy}^{k}$ represents the cumulative mortality rate $k$ days after day $d$, $PM2.5_{cdmy}$ represents the level of $PM2.5$ at day $d$ in county $c$, $X_{cdmy}$ is a vector of control variables that include precipitation, wind speed, and temperature, $\alpha_{c}$ represents county fixed effects, $\alpha_{sm}$ represents state-by-month fixed effects, $\alpha_{my}$ month-by-year fixed effects, and $\epsilon_{cdmy}$ is the disturbance term.
For identification, the authors use wind direction changes as an instrument for air pollution. The authors assume that the first-stage regression presents a parametric nonlinear specification.
$PM2.5_{cdmy}=\sum_{g=1}^{50}f^{g}(\theta_{cdmy})+X_{cdmy}^{\prime}\delta+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (7)
where:
$f^{g}(\theta_{cd})=\gamma_{g}^{1}\mathbf{1}\{G_{c}=g\}sin(\theta_{cd})+\gamma_{g}^{2}\mathbf{1}\{G_{c}=g\}sin(\theta_{cd}/2).$ (8)
In this equation, $\mathbf{1}\{G_{c}=g\}$ is an indicator function that is equal to 1 if county c is member of group $g$ and 0 otherwise. $\theta_{cd}$ is local wind direction measured in radians. There are 100 instrumental variables.
The authors show that their results are robust to different parametric specifications of the first-stage regression but do not allow for a nonparametric specification. However, the concern is that the parametric specifications do not fully capture the nonlinearity in the first-stage equation. Thus, we propose to use a super learner to estimate the nonlinear first-stage regression as follows:
$PM2.5_{cdmy}=h(\theta_{cdmy}),X_{cdmy}^{\prime})+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (9)
where $h(\cdot)$ is an unknown nonlinear function.
4 The Setup
We study a model composed of a linear structural equation, the outcome variable ($y_{it}$) explained by an endogenous variable ($x_{1it}$), $K$ exogenous regressors ($\tilde{x}_{it}$), and an additive individual-specific effect ($\alpha_{i,y}$). The model presents an external instrumental variable ($z_{it}$) for the endogenous regressor. The endogenous regressor is mean dependent on the exogenous regressors, the instrumental variable, and the individual specific effects $\alpha_{i,x}$. We assume that:
Assumption 1.
We observe a sequence of data sets $\{y_{i},x_{1i},...,x_{Ki},z_{i}\}_{i=1}^{N}$, which are $T\times 1$ identical and independent copies of the $K+2$ random vectors $(\bm{y},\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{K},\bm{z})$ taking values in $[-M_{y},M_{y}]^{T}\times[-M_{x1},M_{x1}]^{T}\times[-M_{xK},M_{xK}]^{T}\times[-M_{z},M_{z}]^{T}$ with $M_{y}<\infty$, $M_{xk}<\infty$, $M_{z}<\infty$. Thus, each element $t$ of $\bm{y}$ is bounded and contained in $[-M_{y},M_{y}]$. Similarly, each element of $x_{k}$ is contained in $[-M_{k},M_{k}]$, which means that $\bm{x}_{kt}$ is a continuously distributed random variable with compact support $[-M_{k},M_{k}]$. In addition, each element of $\bm{z}$ is a continuously distributed random variable with compact support $[-M_{z},M_{z}]$.
Assumption 2.
The structural equation has a linear form with an additive unobserved error term and an additive individual specific effect $\alpha_{i,y}$
$\begin{gathered}y_{it}=x_{1it}\beta_{1_{o}}+\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (10)
where $\tilde{x}_{it}=[x_{2it},x_{3it},...,x_{Kit}]^{\prime}$.
The unobserved random term $\varepsilon_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,y}$, and $Var(\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$. The unobserved parameter of interest is $\beta_{1_{o}}$.
Under assumption 2, $\tilde{x}_{it}$ are strictly exogenous in the structural equation conditional on $\alpha_{i,y}$. In addition, $\{z_{is},\quad s=1,...,T\}$ are uncorrelated with the structural error term $\varepsilon_{it}$ conditional on $\alpha_{i,y}$.
Assumption 3.
The regressor $x_{1it}$ is endogenous and has an unknown nonlinear reduced form equation with an additive separable individual specific effect $\alpha_{i,x}$:
$\begin{gathered}x_{1it}=g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,1x}+u_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x}]=0.\end{gathered}$ (11)
with the unknown function $g_{o}:[-M_{x},M_{x}]^{K}\times[-M_{z},M_{z}]\rightarrow\mathcal{G}\subset\mathbb{R}$.
The reduced form error term $u_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,x}$, and $Var(u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x})<\infty$.
Under assumption 3, $\tilde{x}_{it}$, $z_{it}$ are strictly exogenous in the reduced form equation conditional on the individual specific effect $\alpha_{i,1x}$. In addition, this assumption states that the endogenous regressor has a nonlinear relationship with the exogenous regressors $\tilde{x}_{it}$, and the variable $z_{it}$.
As a consequence of Assumptions 2 and 3, $z_{it}$ is an instrumental variable.
The decomposition of $x_{1it}$ into a conditional expectation and an additive disturbance term is appropriate under the assumption that $x_{1it}$ is continuous (Assumption 1). Assumption 3 needs to be modified when $x_{1it}$ is a non-continuous endogenous explanatory variable.
Assumption 4.
The structural error term $\varepsilon_{it}$ has a linear relationship with $u_{it}$:
$\begin{gathered}\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[\omega_{it}|u_{i},\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (12)
Assumption 4 states that $\omega_{it}$ is strictly exogenous of $u_{it}$. The error term $\omega_{it}$ is is independent and identically distributed with a continuous distribution. It has a bounded compact support with zero mean conditional on $u_{i}=[u_{i1},u_{i2},,...,u_{iT}]^{\prime}$, $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,y}$, and $Var(\omega_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$.
As consequence of Assumptions 3 and 4, the structural error term $\varepsilon_{it}$ is correlated with $x_{1it}$. The reason is that $E[x_{1it}\varepsilon_{it}]=E[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})\varepsilon_{it}]$ by Assumption 3. The last expression is equal to $E[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})(\rho_{o}u_{it}+\omega_{it})]=\rho_{o}E[u_{it}^{2}]$ by Assumption 4.
The error term $\omega_{it}$ has zero mean conditional on $x_{1it}$, $\tilde{x}_{it}$, $\alpha_{i,y}$, and $u_{it}$ ($E[\omega_{it}|x_{1it},\tilde{x}_{it},\alpha_{i,y}]=0$) because $E[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y}]=E[\omega_{it}|\tilde{x}_{it},z_{it},\alpha_{i,y},u_{it}]=0$ where the first equality holds because $x_{1it}$ is a one-to-one function of $z_{it}$ conditional on $\tilde{x}_{it}$, $u_{it}$, and $\alpha_{i,y}$ (Wooldridge, 2010). The second equality holds by Assumption 4.
5 Identification: Control Function Approach
The identification of the parameter of interest $\beta_{1_{o}}$ relies on the presence of an instrumental variable $z_{it}$ conditioning on the individual-specific effects (Assumptions 2 and 3). Since we run into an incidental parameter problem, we get rid of the individual-specific effects by using an appropriate transformation of the data $\mathbf{T}:\mathbb{R}\rightarrow\mathbb{R}$. The typical transformations of the data are first-differencing and within transformation. The first-difference transformation is equal to $y_{it}-y_{it-1}$, and the Within transformation is equal to $y_{it}-\sum_{t=1}^{T}y_{it}/T$.
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\},$ (13)
$\mathbf{T}x_{1it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})+\mathbf{T}u_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (14)
where $t_{a}$ is the first time period available after transforming the data. In the case of first-differencing $t_{a}=2$, and for a Within-transformation $t_{a}=1$.
Now, it is tempting to replace equation 14 into 13 to obtain:
$\mathbf{T}y_{it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (15)
In equation 15, one can believe that the estimation of the unknown parameters is straightforward since the error term $\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it}$ in 15 has zero mean conditional $\tilde{x}_{it-1}$, $\tilde{x}_{it-1}$, $z_{it}$, and $z_{it-1}$ as a consequence of assumptions 2, and 3. Moreover, one can think that the best estimation procedure is to estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ in the first stage and in the second stage, plug it in equation 15 and perform simple OLS. This is not possible because the regularization bias of the first stage estimator of $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ contaminates the estimation of $\beta_{1_{o}}$ and $\beta_{2_{o}}$ (Chernozhukov et al., 2018; Robinson, 1988; Guo et al., 2022) and produces estimates that are not $\sqrt{N}$-consistent.
In order to deal with this problem, we can use a control function approach and augment the structural equation by controlling for the unobservable $\mathbf{T}u_{it}$ as follows:
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (16)
In model 16, we can guarantee zero correlation between $\mathbf{T}x_{1it}$, and $\mathbf{T}u_{it}$ thanks to the presence of an instrumental variable in the reduced form equation. In addition, we can set up the following moment conditions:
$E[\phi(\mathbf{T}H_{it};\theta_{o},\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))]=\textbf{0},$ (17)
where $\phi(\mathbf{T}H_{it};\theta_{o},\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))$ is a vector of score functions equal to $\mathbf{T}\omega_{it}\mathbf{T}H_{it}$, $\theta_{o}=[\beta_{1_{o}}\quad\beta_{2_{o}}^{\prime}\quad\rho_{o}]^{\prime}$, $\mathbf{T}\omega_{it}=\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1_{o}}-\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}-\rho_{o}\mathbf{T}u_{it}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, and $\mathbf{T}H_{it}=[\mathbf{T}x_{1it}\quad\mathbf{T}\tilde{x}_{it}\quad\mathbf{T}u_{it}]^{\prime}.$
These moment conditions can be seen as the first-order conditions of the minimization of the following loss function:
$Q_{o}(\mathbf{T}H_{it};\theta,\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))=E[(\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1}-\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k}-\mathbf{T}u_{it}\rho)^{2}].$
This interpretation is important for consistency of the estimator proposed in Section 6 because it permits to specify the necessary primitive condition that $Q_{o}(\mathbf{T}H_{it};\theta,\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))$ has a unique global optimizer $\theta_{o}$ (Newey and McFadden, 1994).
Finally, we identify $\mathbf{T}u_{it}$ since it has zero mean conditional on an appropriate information set $I_{t}=\{\tilde{x}_{it-jk},z_{it-j},0\leq j\leq A\}$ with $A\leq t-1$ as a consequence of Assumption 3.
Example.
First-difference transformation: We could follow Baltagi et al. (2002) by first-differencing the structural and reduced-form equations such that we obtain:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\Delta\varepsilon_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (18)
$\Delta x_{1it}=\Delta g_{o}(\tilde{x}_{it},z_{it})+\Delta u_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (19)
Now, we can use the proposed control function approach and augment the structural equation by controlling for the unobservable $\Delta u_{it}$ as follows:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (20)
Next, we set up the following moment conditions:
$E[\Delta\omega_{it}\Delta H_{it}]=\textbf{0},$ (21)
where $\Delta\omega_{it}=\Delta y_{it}-\Delta x_{1it}\beta_{1_{o}}-\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}-\rho_{o}\Delta u_{it}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, and $\Delta H_{it}=[\Delta x_{1it}\quad\Delta\tilde{x}_{it}\quad\Delta u_{it}]^{\prime}.$
Finally, we identify $\Delta u_{it}$ since it has zero mean conditional on $\tilde{x}_{it}$, $\tilde{x}_{it-1}$, $z_{it}$, $z_{it-1}$ as a consequence of Assumption 3.
Proposition 1.
Under assumption 3, $\Delta g({\tilde{x}_{it},z_{it}})$ is identified by $E[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$.
Proof 1.
$E[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=E[\Delta g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]+E[\Delta u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$ (22)
By the linearity property of the expectation operator, we have that the first term is equal to:
$\begin{split}E[g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-E[g(\tilde{x}_{it-1},z_{it-1})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\\ g(\tilde{x}_{it},z_{it})-g(\tilde{x}_{it-1},z_{it-1})\end{split}$ (23)
Similarly for the second term, by the Law of Iterated Expectations and Assumption 3:
$E[u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-E[u_{it-1}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=0$ (24)
Remark 1.
A similar analysis and conclusion is obtained using a Within transformation which is equivalent to time demeaning the data. In this case, the information set $I_{t}=\{\tilde{x}_{it-j},z_{it-j},0\leq j\leq t-1\}$.
In Section 6, we propose an estimation method based on the Control Function approach.
6 Super Learner Control Function estimation
As described in the previous section, we can write the transformed structural equation for observation $i$ at period $t$ as follows:
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\mathbf{T}\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (25)
In this transformed equation, the parameters of interest are identified because $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$ and $\mathbf{T}u_{it}$ are uncorrelated with error term $\mathbf{T}\omega_{it}$ under assumption 4.
In addition, it is clear that if $\mathbf{T}u_{it}$ is observed, we could estimate the three parameters $\beta_{1_{o}}$, $\beta_{2_{o}}$, $\rho_{o}$ using the sample counterparts of the population moment conditions 17. But since $\mathbf{T}u_{it}$ is not observed, we propose the following estimation procedure:
First Step:
We partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$. This sample splitting is equal to the one proposed by Emmenegger and Bühlmann (2023) and it was developped without knowledge of it. Alternatively, if the cluster structure in the data is known (individuals exhibit correlation within clusters but not across them, and both cluster membership and the number of clusters are identified) the data should be partitioned accordingly. Then, we estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})=E[\mathbf{T}x_{1it}|I_{t}]$ using a super learner with partition $S_{b}^{c}=\{i\in S_{g},g\neq b\}$. Then, we obtain the residuals $\widehat{\mathbf{T}u}_{it}=x_{1it}-\widehat{\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}=\{i\in S_{b}\}$.
Finally, the estimator of $\theta_{o}$ for partition $S_{b}$ is the solution of the following sample moment conditions:
$\frac{1}{|S_{b}|(T-t_{a})}\sum_{i\in S_{b}}\sum_{t=t_{a}}^{T}\phi(\widehat{\mathbf{T}H}_{it};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it}))=0,$ (26)
where $\widehat{\mathbf{T}H}_{it}=[\mathbf{T}x_{1it}\quad\mathbf{T}\tilde{x}_{it}\quad\widehat{\mathbf{T}u}_{it}]^{\prime}$, and $\widehat{\mathbf{T}u}_{it}$ are the residuals obtained in the first stage for partition $S_{b}$.
This is equivalent to estimate $\theta_{o}$ by performing ordinary least squares regression of $\mathbf{T}y_{it}$ on $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$ and $\widehat{\mathbf{T}u}_{it}$ or minimizing the following quadratic loss function :
$\hat{\theta}=\begin{subarray}{c}argmin\\ \theta\end{subarray}Q_{b}(\widehat{\mathbf{T}H};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it})),$
with $Q_{b}(\widehat{\mathbf{T}H};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it}))=\frac{1}{|S_{b}|(T-t_{a})}\sum_{i\in S_{b}}\sum_{t=t_{a}}^{T}(\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1}-\mathbf{T}\tilde{x}_{it}^{\prime}\beta_{2}-\widehat{\mathbf{T}u_{it}}\rho)^{2}$, $\tilde{X}$ a $N(T-t_{a})\times(K-1)$ vector collecting all observations of $\tilde{x}_{it}$, and $\widehat{\mathbf{T}H}$ a $|S_{b}|(T-t_{a})\times(K+1)$ matrix collecting all observations of $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$, and $\widehat{\mathbf{T}u}_{it}$. Importantly, ordinary least squares estimation is not the most efficient if $T>2$, or if the $\varepsilon_{it}$ is not homoskedastic neither uncorrelated across both $i$ and $t$. In this situation, a most efficient estimation is Generalized Least Squares estimation or another appropriate estimation method that accounts for this.
Second Step:
We average $\hat{\theta}_{b}$ to obtain the final estimator of $\theta_{o}$ as:
$\hat{\theta}_{o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\theta}_{b}.$
Thus, the estimator of our parameter of interest $\beta_{1o}$ is:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\beta}_{b}.$
Algorithm 1 Super Learner Control Function estimation
  Input: Panel data sets $\{y_{i},X_{i},z_{i}\}_{i\in\{1,2,...,N\}}$ from model satisfying assumptions 1 to 4, a natural number B.
  Output: An estimator of the parameter of interest $\beta_{1o}$.
  Split the individual index set $\{1,2,...,N\}$ into B sets $S_{1},S_{2},...,S_{B}$.
  for $b\in\{1,2,...,B\}$ do
    Estimate the conditional expectation $E[\mathbf{T}x_{1it}|I_{t}]$ using a super learner and data correspondig to $S_{b}^{c}$.
    Predict $E[\mathbf{T}x_{1it}|I_{t}]$ for $i\in S_{b}$.
    Obtain the residuals $u_{it}$ for $i\in S_{b}$ using the prediction of $E[\mathbf{T}x_{1it}|I_{t}]$ of previous step.
    Estimate $\theta_{o}=[\beta_{1o},\beta_{2o},\rho_{o}]$ solving the sample moment conditions 26 for $i\in S_{b}$.
  end for
  Compute $\hat{\theta}_{o}=\frac{\sum_{1}^{B}\widehat{\theta_{b}}}{B}.$
Example.
First-difference transformation (continuation):
As described in the previous section, we can write the first-differenced structural equation for observation $i$ at period $t$ as follows:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (27)
First, we partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and we estimate $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner with partition $S_{b}^{c}$. Then, we obtain the residuals $\widehat{\Delta u}_{it}=\Delta x_{1it}-\widehat{\Delta g_{o}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}$. We obtain the estimator of $\theta_{o}$ is the solution of the following sample moment conditions:
$\frac{1}{|S_{b}|(T-1)}\sum_{i\in S_{b}}\sum_{t=2}^{T}\phi(\widehat{\Delta H}_{it};\theta,\widehat{\Delta g}_{o}(\tilde{x}_{it},z_{it}))=0,$ (28)
where $\widehat{\Delta H}_{it}=[\Delta x_{1it}\quad\Delta\tilde{x}_{it}\quad\widehat{\Delta u}_{it}]^{\prime}$, and $\widehat{\Delta u}_{it}$ are the residuals obtained in the first stage for partition $S_{b}$.
Finally, we estimate our parameter of interest $\beta_{1o}$ using:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\beta}_{b}.$
Algorithm 2 Super Learner Control Function Estimation: First-difference transformation
  Input: Panel data sets $\{y_{i},X_{i},z_{i}\}_{i\in\{1,2,...,N\}}$ from model satisfying assumptions 1 to 4, a natural number B.
  Output: An estimator of the parameter of interest $\beta_{1o}$.
  Split the individual index set $\{1,2,...,N\}$ into B sets $S_{1},S_{2},...,S_{B}$.
  for $b\in\{1,2,...,B\}$ do
    Estimate the conditional expectation $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner and data correspondig to $I_{b}^{c}$.
    Predict $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ for $i\in S_{b}$.
    Obtain the residuals $u_{it}$ for $i\in I_{b}$ using the prediction of $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ of previous step.
    Estimate $\theta_{o}=[\beta_{1o},\beta_{2o},\rho_{o}]$ solving the sample moment conditions 26 for $i\in S_{b}$.
  end for
  Compute $\bar{\hat{\beta}}_{1o}=\frac{\sum_{1}^{B}\widehat{\beta_{1o,b}}}{B}.$
7 Monte Carlo simulation
We test the performance of the proposed estimation methods using a Monte Carlo simulation experiment. As competing methods, we consider Within OLS, and Within 2SLS estimators including polynomial transformations of the covariates and instruments of degree 1 and 5.
7.1 Settings
7.1.1 Linear structural equation and non-linear secondary equation
We generate 100 samples from a triangular simultaneous model with a linear structural equation and a non-linear reduced-form equation with $N=1000$ and $T=2$. Both equations present additive disturbance terms, and individual-specific effects. The structural equation presents one endogenous regressor $x_{1it}$ and one exogenous covariate $x_{2it}$.
$\begin{split}y_{it}=\beta_{1}x_{1it}+\beta_{2}x_{2it}+\alpha_{i}+\epsilon_{it},\end{split}$ (29)
where:
$x_{2it}=\alpha_{i}+\zeta_{it},\quad with\quad\zeta_{it}\sim U(-2,2),$ (30)
$z_{it}=\alpha_{i}+\nu_{it},\quad with\quad\nu_{it}\sim U(-2,2),$ (31)
$x_{1it}=g(x_{2it},z_{it})+\alpha_{i}+u_{it},$ (32)
$u_{it}\sim U(-1,1).$ (33)
$\epsilon_{it}=0.9*u_{it}+\tilde{\zeta_{it}},\quad with\quad\tilde{\zeta}_{it}\sim U(-1,1).$ (34)
$\alpha_{i}\sim U(-1,1).$ (35)
The unknown function for $g(x_{2it},z_{it})$ is similar to the ones proposed by Guo et al. (2022), and depends on the parameter $a$ that controls the nonlinearity in the functional form and the strength of the linear correlation between $x_{1it}$ and $z_{it}$. As in Guo et al. (2022), a larger a causes higher nonlinearity and lower linear correlation between $x_{1it}$ and $z_{it}$. However, we cannot compare our results to theirs because they consider a model for cross-sectional data.
Table 1: Monte Carlo Experiment
Scenario $g(\cdot)$ $\beta_{1}$ $dim(x_{2})$ $\beta_{2}$
1 $-a\cdot|z_{it}|-2tanh(x_{2it})+1/a*z_{it}$ 1 1 1
((a))
((b))
Figure 1: a) Relationship between $\Delta x_{1it}$ and $z_{it}$ with $a=10$, b) Relationship between $\Delta x_{1it}$ and $z_{it}$ with $a=1$
7.2 The results
In this section, we present the results of the simulation experiment. In figure 2, we present the average estimated parameter of interest $\beta_{1o}$ for different values of $a$. The methods are Within OLS (WOLS), Within 2SLS (W2SLS) using polynomial of degree 1 of the covariate and the instrumental variables, Within 2SLS using polynomial of degree 5 of the covariate and the instrumental variables (W2SLS_poly), the Super Learner Control Function estimator on the First-Differenced data (FDSLCF), and the Super Learner Control Function estimator on the Within-transformed data (WSLCF). In addition, we plot the average correlation between the endogenous covariate and the instrumental variable. We also plot the p-value of the estimated control functions using the first-differenced and the within transformed data. As $a$ is larger, the nonlinearity in the reduced form equation is more important. This translates in a lower correlation between the endogenous regressor and the instrumental variables. As a result, the Within-2SLS estimators presented a decay in their performance. In contrast, the Super Learner Control Function estimators remain stable around the true value of the parameter of interest $\beta_{1o}=1$.
In panel (a) of figure 3, we present the boxplot of the estimated parameter of interest $\beta_{1o}$ for a highly nonlinear reduced form equation (DGP 1 with $a=5$), and in panel (b) for a linear reduced form equation (DGP 1 with $a=1$). The estimated parameters with the Super Learner Control Function estimators are tightly concentrated around the true value of the parameter of interest. In contrast, the Within estimates present wide variation around the true value.
For the estimation of the nuisance parameters, we used a super learner with individual learners the mean, the linear regression, and a feed-forward neural network with one hidden layer.
Figure 2: Average estimates of $\beta_{1o}$ for 100 samples simulated with different values of $a$, $N=1000$, $T=2$
((a))
((b))
Figure 3: a) Estimates of $\beta_{1o}$ for a DGP highly nonlinear (a=5), b) Estimates of $\beta_{1o}$ for a DGP highly linear (a=1), $N=1000$, $T=2$
8 Empirical Application: Air Pollution and Educational Outcomes in the U.S. (Revisited)
As an illustration, we estimate the causal effect of air pollution on student performance in the United States. The data on air pollution are obtained from Deryugina et al. (2019). This dataset presents daily data on air pollutants: PM 2.5, ozone ($O_{3}$), carbon monoxide (CO), sulfur dioxide ($SO_{2}$), nitrogen dioxide ($NO_{2}$). It also contains data on wind direction, wind speed, and weather conditions at county level. We aggregate the data to yearly level by averaging across days. The data on student performance are obtained from Stanford Educational Data Archive (SEDA) Fahle et al. (2024) for county-yearly level. We merge both datasets by county-year indicator, and keep cohort 2006. Our final sample is composed of years 2009 to 2013 and 3065 counties.
We specify the following linear regression for subject district $d$, at year $t$:
$y_{dt}=\beta PM2.5_{dt}+X_{dt}^{\prime}\gamma+\alpha_{d}+\lambda_{t}+\epsilon_{dt},$ (36)
where $y_{dt}$ represents the average grade of subject mathematics, county $d$ at year $t$. $PM2.5_{dt}$ represents the yearly average of $PM2.5$ concentration at county $d$ in year $t$, $X_{dt}$ is a vector of control variables at county-year level including average school characteristics, social characteristics, and weather variables. $\alpha_{d}$ represents county fixed effects, and $\lambda_{t}$ represents time fixed effects.
As mentioned before, $PM2.5_{dt}$ is endogenous. Thus, following Deryugina et al. (2019) we use wind direction as an instrument of air pollution along with a nonlinear first-stage equation as follows:
$PM25_{dt}=h(WD_{dt},X_{ft})+\eta_{d}+\psi_{t}+\varepsilon_{dt},$ (37)
with $WD_{dt}$ is the yearly average wind direction in county $d$ in year $t$, $\eta_{d}$ are county specific effects, and $\psi_{t}$ are time fixed effects.
We estimate that an increase in the concentration of PM 2.5 causes a reduction in test scores of 0.0151 standard deviations.
Table 2: Estimated effect of air pollution (PM 2.5 concentration) on student performance
WOLS W2SLS FD SLCF
-0.0007 -0.0284 -0.0151
(0.8070) (0.0008) (0.0516)
Note: US counties, 2009-2014 (N=3065, T=6). Standard errors in parenthesis, Bootstrap standard errors for FD SLCF, W SLCF.
9 Conclusions
This paper proposes a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation to deal with the issue of weak instrumental variables due to nonlinearities in panel data settings. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. We estimate the control function using a super learner using sample splitting. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the estimators perform well, provided that we can learn accurately the nuisance parameter in the first stage.