### v6

# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments (instruments are weakly correlated with the endogeneous covariate). As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. The estimator is consistent and asymptotically normal achieving a parametric rate of convergence. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators. Finally, we show that the SLCF estimator is not numerically equivalent to a naive plugin IV estimator.
Keywords— Endogeneity, Super Learner, Instrumental Variables, Identification, Control Function, Panel data, Individual Fixed Effects
1 Introduction
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders, and some of them are time-invariant. In this setup, it is common to assume that the structural is linear and to use a linear projection in the reduced form equation. However, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to problems due to weak instruments since the strength of the linear relationship might not be high. This paper proposes to exploit the nonlinearties in the relationship between the endogenous covariate and the instrumental variables as a way to avoid problems of weak instruments due to nonlinearities.  [^1] Instruments are weak when they are weakly correlated to the endogenous covariate. In a cross-sectional setup, weak IVs causes that the median of the 2SLS estimator is biased towards the OLS estimator, 2SLS t-test presents size inflation (Keane and Neal (2024)). In panel data, Within-2SLS estimator is consistent when the time dimension grows (Cai et al., 2012).
An example is the relationship between air pollution and children’s educational outcomes, where both variables are influenced by unobserved confounders that can be either time-varying or time-invariant. For instance, neighborhood characteristics and family income are often unobserved in publicly available data. While neighborhood characteristics typically remain constant over time, family income fluctuates. Both factors are relevant to a child’s academic achievement and exposure to air pollution, leading to endogeneity in pollution exposure. To address this, fixed effects can be used to control for unobserved neighborhood characteristics, while an instrumental variable can help account for time-varying confounders. A commonly used instrumental variable is wind direction, which has a well-established nonlinear relationship with air pollution levels (Zabrocki et al., 2022).
This paper studies a triangular structural panel data model in which the reduced form equation is nonlinear and individual specific effects are present. It proposes the use of Super Learning for the estimation of the nonlinearities in the model. For simplicity consider the following triangular model without covariates for $i\in\{1,...,N\}$, $t\in\{1,...,T\}$:
$y_{it}=x_{1it}\beta_{1_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad with\quad\mathbb{E}[\varepsilon_{it}|z_{i},\alpha_{i,y}]=0,$ (1)
$x_{1it}=g_{o}(z_{it})+\alpha_{i,1x}+u_{it},\quad with\quad\mathbb{E}[u_{it}|z_{i},\alpha_{i,1x}]=0,$ (2)
$\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad with\quad\mathbb{E}[\omega_{it}|u_{i},z_{i},\alpha_{i,y}]=0,$ (3)
[…]
The model comprises a structural equation (1) with an endogenous regressor $x_{1it}$ that presents a non-linear relationship with the available instrumental variable $z_{it}$ (2). Both equations present individual effects that represent time-invariant unobserved confounders. In this setup, conditioning on the individual effects allows us to control for time-invariant unobserved confounders. But, estimating all the individual effects leads to incidental parameter bias as the number of parameters to estimate grows with the sample size. Then, in order to eliminate the individual effects, we propose to transform the model (e.g. first-differencing, or a within transformation which is equivalent to obtain the deviations of the original variables from their individual mean). On the other hand, the presence of time-varying unobserved confounders is modeled through a linear relationship between the error terms in both equations (3). We deal with this issue by exploiting available exclusion restrictions (instrumental variables) along with a Control Function approach on the transformed model.
[…]
Our paper contributes to the literature on the use of machine learning to estimate panel data models. We are the first to use super learning to estimate a panel data model with an endogenous covariate (due to unobserved confounders or measurement error) and additive individual fixed effects using a control function approach and exploiting nonlinearities. In addition, this paper is the first to use either First-Difference or Within transformation to take care of the individual specific effects while using machine learning methods. In contrast, these transformations have been used in the nonparametric panel data literature. The available papers using machine learning to estimate panel data models focus on exogenous covariates (time-varying confounders are observed) and the presence of individual fixed effects. Klosin and Vilgalys (2022) proposes a debiased LASSO estimator to estimate heterogeneous treatment effects of a continuous treatment, which is modeled using a structural nonparametric panel data model with additive fixed effects and strictly exogenous covariates. They account for the individual specific effects by taking the first-difference of the original model. Semenova et al. (2023) focuses on a structural equation model for panel data allowing for heterogenous treatment effects with unconfoundedness on observables (sequentially exogenous covariates) and individual specific fixed effects. In order to deal with the individual fixed effects, the authors use a Mundlak approach. They provide an algorithm composed of three steps. In the first step, using cross-fitting they estimate the nuisance functions of the structural and treatment equations to obtain the residuals. In the second step, they estimate the conditional average treatment effect using lasso estimators. In the third step, they provide debiased inference.
This paper also contributes to the literature on weak instruments in panel data by being the first one to consider the problem of weak instruments due to nonlinearities, and to use super learning with control function approach as a solution. The available papers on weak instrumental variables in panel data focus on the asymptotic properties of the Within 2SLS estimator showing that it is consistent even with weak instruments when the time dimension grows to infinity. Cai et al. (2012) show that the Within 2SLS is consistent when the individual and time dimension grow to infinity. They also present asymptotical pivotal tests (Anderson-Rubin test, Kleibergen test) for longitudinal data. Cai et al. (2012) provide conditions for consistency of the Within 2SLS estimator.
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5 we describe the identification strategy. In Section 6 we present the SLCF estimator. In Section 8 we present the large sample properties of the SLCF estimator, in Section 10, we present the simulation experiment and its results, in Section 11 we present the results of the empirical application, and in Section 12 we present the conclusions of the paper.
2 Other Related literature
This paper is related to the literature on semiparametric panel data models. Rodriguez-Poo and Soberon (2017) explains that one can either use a first-differencing approach or a profiling technique to estimate a semi-parametric model with an additive separable individual specific effect and a disturbance term. Li (1996) explains that kernel estimation of the conditional expectation of the first-differenced endogenous regressor is not feasible when the dimension of the conditioning set is larger than 5. As a solution, Baltagi et al. (2002) presents an additive estimation of the first-differenced non-linear function.
In addition, this paper is related to the literature on orthogonal/debiased estimation, which focuses on providing $\sqrt{N}$ consistent estimators of low-dimensional parameters in the presence of high-dimensional nuisance parameters. Chernozhukov et al. (2018) defines an orthogonal score as the one that presents a vanishing Gateaux derivative with respect to the nuisance parameters when evaluated at the true finite-dimensional parameter values. The orthogonal score is closely related to the score proposed by Robinson (1988).
Finally, our work relates to the literature of mixed effects models and machine learning. Our model is related to a mixed effects model assuming that covariates are exogenous conditioning on the cluster random effects. Emmenegger and Bühlmann (2023) uses double machine learning to estimate a partially linear mixed-effects model. Young and Shah (2024) proposes a sandwich boosting for accurately estimating partially linear models for grouped data.
2.1 Super Learner
The Super Learner was proposed by Van der Laan et al. (2007), it is a heterogeneous ensemble method within Supervised Learning. It consists of different base learners that are trained on the training data set, then their predictions are combined using a meta-learner by minimizing the cross-validated risk. The meta-learner can be a linear regression of the predictions of each learner on the outcome variable. Van der Laan et al. (2007) shows that the Super Learner performs asymptotically as well as or better than learner in the library of base learners.
3 Motivating examples
As mentioned in the Introduction, measuring the causal effect of air pollution on health and educational outcomes is not straightforward as it is not randomly assigned. More specifically, time-varying unobservables such as family income might affect air pollution and health or educational outcomes. In addition, unobserved time-invariant variables, such as neighborhood characteristics or genetics, might also be present. While time-invariant unobservables could be dealt with by including individual fixed effects in the model, the presence of time-varying unobservable confounders requires instrumental variables. In this setup, wind direction is a commonly used instrumental variable. While it is well documented that the relationship between air pollution and wind direction is nonlinear (Zabrocki et al., 2022), it is common to see that the nonlinearity is ignored or modeled with a parametric assumption.
3.1 Air pollution and educational outcomes
Gilraine and Zheng (2022) studies the effect of air pollution on educational outcomes. They specify the following linear regression for subject $s$, cohort $c$, district $d$, year $t$:
$y_{sdct}=\beta PM2.5_{dt}+X_{sdct}^{\prime}\gamma+W_{dt}^{\prime}\eta+\omega_{s}+\alpha_{d}+\theta_{c}+\lambda_{t}+\epsilon_{scdt},$ (4)
where $y_{sdct}$ represents the average grade of subject $s$, cohort $c$, district $d$ at year $t$. $PM2.5_{dt}$ represents the average level of $PM2.5$ at district $d$ in year $t$, $X_{sdct}$ is a vector of control variables, $W_{dt}$ is a vector of weather variables, $\omega_{s}$ represents subject specific effects, $\alpha_{d}$ represents district fixed effects, $\theta_{c}$ represents cohort fixed effects, $\lambda_{t}$ are time fixed effects, and $\epsilon_{scdt}$ is the disturbance term.
As mentioned before, $PM2.5_{dt}$ is not randomly assigned and subject to measurement error. Thus, the authors use year-to-year changes in coal production and a shift-share instrument equal to the interaction of fuel shares used for nearby production with national growth rates along with a linear first-stage regression.
As an alternative, following Deryugina et al. (2019) we could use wind direction as an instrument of air pollution along with a nonlinear first-stage equation as follows:
$PM25_{dt}=h(WD_{dt},X_{dt})+\gamma_{d}+\varepsilon_{dt},$ (5)
where $WD_{dt}$ is the average wind direction in district $d$ in year $t$, and $\gamma_{d}$ represent district specific effects.
3.2 Air pollution and health outcomes
Deryugina and Reif (2023) studies the effect of air pollution on mortality rates. More precisely, they specify the following linear regression for county $c$, day $d$, month $m$ and year $y$:
$y_{cdmy}^{k}=\beta PM2.5_{cdmy}+X_{cdmy}^{\prime}\gamma+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\epsilon_{cdmy},$ (6)
where $y_{cdmy}^{k}$ represents the cumulative mortality rate $k$ days after day $d$, $PM2.5_{cdmy}$ represents the level of $PM2.5$ at day $d$ in county $c$, $X_{cdmy}$ is a vector of control variables that include precipitation, wind speed, and temperature, $\alpha_{c}$ represents county fixed effects, $\alpha_{sm}$ represents state-by-month fixed effects, $\alpha_{my}$ month-by-year fixed effects, and $\epsilon_{cdmy}$ is the disturbance term.
For identification, the authors use wind direction changes as an instrument for air pollution. The authors assume that the first-stage regression presents a parametric nonlinear specification.
$PM2.5_{cdmy}=\sum_{g=1}^{50}f^{g}(\theta_{cdmy})+X_{cdmy}^{\prime}\delta+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (7)
where:
$f^{g}(\theta_{cd})=\gamma_{g}^{1}\mathbf{1}\{G_{c}=g\}sin(\theta_{cd})+\gamma_{g}^{2}\mathbf{1}\{G_{c}=g\}sin(\theta_{cd}/2).$ (8)
In this equation, $\mathbf{1}\{G_{c}=g\}$ is an indicator function that is equal to 1 if county c is member of group $g$ and 0 otherwise. $\theta_{cd}$ is local wind direction measured in radians. There are 100 instrumental variables.
The authors show that their results are robust to different parametric specifications of the first-stage regression but do not allow for a nonparametric specification. However, the concern is that the parametric specifications do not fully capture the nonlinearity in the first-stage equation. Thus, we propose to use a super learner to estimate the nonlinear first-stage regression as follows:
$PM2.5_{cdmy}=h(\theta_{cdmy},X_{cdmy}^{\prime})+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (9)
where $h(\cdot)$ is an unknown nonlinear function.
4 The Setup
We study a model composed of a linear structural equation, the outcome variable ($y_{it}$) explained by an endogenous variable ($x_{1it}$), K1K-1 exogenous regressors ($\tilde{x}_{it}$), and an additive individual-specific effect ($\alpha_{i,y}$). The model presents an external instrumental variable ($z_{it}$) for the endogenous regressor. The endogenous regressor is mean dependent on the exogenous regressors, the instrumental variable, and the individual specific effects $\alpha_{i,x}$. We assume that:
Assumption 1.
We observe a sequence of data sets $\{O_{i}=(y_{i},x_{1i},...,x_{Ki},z_{i})\}_{i=1}^{N}$, which are $T_{i}\times 1$ independent copies of the $K+2$ random vectors $(\mathbf{y},\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{K},\mathbf{z})$ taking values in $M_{y},M_{y}]^{T}\times[-M_{x1},M_{x1}]^{T}\times[-M_{xK},M_{xK}]^{T}\times[-M_{z},M_{z}]^{T}$ with $M_{y}<\infty$, $M_{xk}<\infty$, $M_{z}<\infty$. Thus, each element $t$ of $\mathbf{y}$ is bounded and contained in $M_{y},M_{y}]$. Similarly, each element of $x_{k}$ is contained in $M_{xk},M_{xk}]$, which means that $\mathbf{x}_{kt}$ is a continuously distributed random variable with compact support $M_{xk},M_{xk}]$. In addition, each element of $\mathbf{z}$ is a continuously distributed random variable with compact support $M_{z},M_{z}]$.
Assumption 1 states that each individual $i$ presents $T_{i}$ observations. We assume that $T_{i}$ is uniformly bounded such that $T_{i}<T_{max}$. The total number of observations is denoted by $N_{T}=\sum_{i}T_{i}$.
Assumption 2.
The structural equation has a linear form with an additive unobserved error term and an additive individual specific effect $\alpha_{i,y}$
$\begin{gathered}y_{it}=x_{1it}\beta_{1_{o}}+\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\alpha_{i,y}+\varepsilon_{it},\qquad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (10)
where $\tilde{x}_{it}=[x_{2it},x_{3it},...,x_{Kit}]^{\prime}$.
The unobserved random term $\varepsilon_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,y}$, and $Var(\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$. The unobserved parameter of interest is $\beta_{1_{o}}$.
Under assumption 2, $\tilde{x}_{it}$ are strictly exogenous in the structural equation conditional on $\alpha_{i,y}$. In addition, $\{z_{is},\quad s=1,...,T\}$ are uncorrelated with the structural error term $\varepsilon_{it}$ conditional on $\alpha_{i,y}$.
Assumption 3.
The regressor $x_{1it}$ is endogenous and has an unknown nonlinear reduced form equation with an additive separable individual specific effect $\alpha_{i,x}$:
$\begin{gathered}x_{1it}=g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,1x}+u_{it},\qquad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x}]=0.\end{gathered}$ (11)
with the unknown function $g_{o}:[-M_{x},M_{x}]^{K}\times[-M_{z},M_{z}]\rightarrow\mathcal{S}\subset\mathbb{R}$. We do not impose any restriction on the form of $g_{o}$.
The reduced form error term $u_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,x}$, and $Var(u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x})<\infty$.
Under assumption 3, $\tilde{x}_{it}$, $z_{it}$ are strictly exogenous in the reduced form equation conditional on the individual specific effect $\alpha_{i,1x}$. In addition, this assumption states that the endogenous regressor has a nonlinear relationship with the exogenous regressors $\tilde{x}_{it}$, and the variable $z_{it}$.
As a consequence of Assumptions 2 and 3, $z_{it}$ is an instrumental variable.
The decomposition of $x_{1it}$ into a conditional expectation and an additive disturbance term is appropriate under the assumption that $x_{1it}$ is continuous (Assumption 1). Assumption 3 needs to be modified when $x_{1it}$ is a non-continuous endogenous explanatory variable.
Assumption 4.
The structural error term $\varepsilon_{it}$ has a linear relationship with $u_{it}$:
$\begin{gathered}\varepsilon_{it}=\rho u_{it}+\omega_{it},\qquad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[\omega_{it}|u_{i},\tilde{x}_{i},z_{i},\alpha_{i,y},\alpha_{i,1x}]=0.\end{gathered}$ (12)
Assumption 4 states that $\omega_{it}$ is strictly exogenous of $u_{it}$. The error term $\omega_{it}$ is is independent and identically distributed with a continuous distribution. It has a bounded compact support with zero mean conditional on $u_{i}=[u_{i1},u_{i2},,...,u_{iT_{i}}]^{\prime}$, $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,y}$, and $Var(\omega_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})=\sigma^{2}_{\omega}<\infty$. This assumption is relaxed in a follow up study.
As consequence of Assumptions 3 and 4, the structural error term $\varepsilon_{it}$ is correlated with $x_{1it}$. The reason is that $\mathbb{E}[x_{1it}\varepsilon_{it}]=\mathbb{E}[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})\varepsilon_{it}]$ by Assumption 3. The last expression is equal to $\mathbb{E}[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})(\rho_{o}u_{it}+\omega_{it})]=\rho_{o}\mathbb{E}[u_{it}^{2}]$ by Assumptions 3 and 4.
The error term $\omega_{it}$ has zero mean conditional on $x_{1it}$, $\tilde{x}_{it}$, $\alpha_{i,y}$, and $u_{it}$ ($\mathbb{E}[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y}]=0$) because $\mathbb{E}[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y},\alpha_{i,1x}]=\mathbb{E}[\omega_{it}|\tilde{x}_{it},z_{it},\alpha_{i,y},\alpha_{i,1x},u_{it}]=0$ where the first equality holds because $x_{1it}$ is a one-to-one function of $z_{it}$ conditional on $\tilde{x}_{it}$, $u_{it}$, and $\alpha_{i,1x}$ (Wooldridge, 2010). The second equality holds by Assumption 4. Then, we can apply the Law of Iterated Expectations to proof the claim.
5 Identification: Control Function Approach
The identification of the parameter of interest $\beta_{1_{o}}$ relies on the presence of an instrumental variable $z_{it}$ conditioning on the individual-specific effects (Assumptions 2 and 3). Since we run into an incidental parameter problem, we get rid of the individual-specific effects by using an appropriate transformation of the data $\tau:\mathbb{R}\rightarrow\mathbb{R}$. The typical transformations of the data are first-differencing and within transformation. The first-difference transformation is equal to $y_{it}-y_{it-1}$, and the Within transformation is equal to $y_{it}-\sum_{t=1}^{T_{i}}y_{it}/T_{i}$.
$\tau y_{it}=\tau x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\tau x_{kit}\beta_{2k_{o}}+\tau\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\},$ (13)
$\tau x_{1it}=\tau g_{o}(\tilde{x}_{it},z_{it})+\tau u_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (14)
where $t_{a}$ is the first time period available after transforming the data. In the case of first-differencing $t_{a}=2$, and for a Within-transformation $t_{a}=1$.
Now, it is tempting to replace equation 14 into 13 to obtain:
$\tau y_{it}=\tau g_{o}(\tilde{x}_{it},z_{it})\beta_{1_{o}}+\sum_{k=2}^{K}\tau x_{kit}\beta_{2k_{o}}+\tau u_{it}\beta_{1_{o}}+\tau\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (15)
In equation 15, one can believe that the estimation of the unknown parameters is straightforward since the error term $\tau u_{it}\beta_{1_{o}}+\tau\varepsilon_{it}$ in 15 has zero mean conditional on an appropriate information set $I_{t}=\{\tilde{x}_{it-jk},z_{it-j},B\leq j\leq D\}$ with , $B\geq T_{i}-1$, $D\leq t-1$ as a consequence of assumptions 2, and 3. Moreover, one can think that the best estimation procedure is to estimate $\tau g_{o}(\tilde{x}_{it},z_{it})$ in the first stage and in the second stage, plug it in equation 15 and perform simple OLS. But this is not possible because the regularization bias of the first stage estimator of $\tau g_{o}(\tilde{x}_{it},z_{it})$ contaminates the estimation of $\beta_{1_{o}}$ and $\beta_{2_{o}}$ (Chernozhukov et al., 2018; Robinson, 1988; Guo et al., 2022) and produces estimates that are not $\sqrt{N_{T}}$-consistent.
In order to deal with this problem, we can use a control function approach and augment the structural equation by controlling for the unobservable $\tau u_{it}$ as follows:
$\tau y_{it}=\tau x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\tau x_{kit}\beta_{2k_{o}}+\rho_{o}\tau u_{it}+\tau\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (16)
In model 16, we can guarantee zero correlation between $\tau x_{1it}$, and $\tau\omega_{it}$ thanks to the presence of an instrumental variable in the reduced form equation.
If we stack up the time observations and denote as $M_{\tau_{i}}$ the transformation operator that maps $\mathbb{R}^{T_{i}}\rightarrow\mathbb{R}^{T_{i,a}}$ with $T_{i,a}$ equal to $T_{i}$ if we use the within transformation and $T_{i}-1$ if we use the first-difference transformation, we set up the following moment conditions:
$\mathbb{E}[\phi(O_{i};\theta_{o},\tau g_{o})]=\textbf{0},$ (17)
where $\phi(O_{i};\theta_{o},\tau g_{o})$ is a vector of score functions equal to $(M_{\tau_{i}}H_{i})^{\prime}V_{i}^{-1}M_{\tau_{i}}\omega_{i}$, $\theta_{o}=[\beta_{1_{o}}\quad\beta_{2_{o}}^{\prime}\quad\rho_{o}]^{\prime}$, $M_{\tau_{i}}\omega_{i}=M_{\tau_{i}}y_{i}-M_{\tau_{i}}x_{1i}\beta_{1_{o}}-\sum_{k=2}^{K}M_{\tau_{i}}x_{ki}\beta_{2k_{o}}-\rho_{o}M_{\tau_{i}}u_{i}$, $\textbf{0}=[0\quad\cdots\quad 0\quad 0]^{\prime}$, and $M_{\tau_{i}}H_{i}=[M_{\tau_{i}}x_{1i}\quad M_{\tau_{i}}\tilde{x}_{i}\quad M_{\tau_{i}}u_{i}]$, and $V_{i}=\mathbb{E}[(M_{\tau_{i}}\omega_{i})(M_{\tau_{i}}\omega_{i})^{\prime}|H_{i},z_{i}]=\sigma^{2}_{\omega}M_{\tau_{i}}M_{\tau_{i}}^{\prime}=\sigma^{2}_{\omega}\tilde{V}_{i}$.
Remark 1.
If we use the first-difference transformation, $M_{\tau_{i}}=D_{i}=\begin{bmatrix}1&-1&0&\cdots&0\\ 0&1&-1&\cdots&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&-1\\ \end{bmatrix}$, with dimensions $(T_{i}-1)\times T_{i}$. If we use the within transformation, $M_{\tau_{i}}=W_{i}=I_{i}-1/T_{i}\begin{bmatrix}1&1&1&\cdots&1\\ 1&1&1&\cdots&1\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 1&\cdots&1&1&1\\ \end{bmatrix}$ , with dimensions $T_{i}\times T_{i}$, $I_{i}$ is the identity matrix.
Remark 2.
The moment condition is valid even under misspecification of $V_{i}$ because we assume strict exogeneity. If we would like to relax the assumption of strict exogeneity to weak exogeneity, then it is better to use $\mathbb{E}[(M_{\tau_{i}}H_{i})^{\prime}M_{\tau_{i}}\omega_{i}]=0$ as moment conditions.
Remark 3.
If $M_{\tau_{i}}=W_{i}$ (the within transformation), the moment condition can be simply written as $\mathbb{E}[(M_{\tau_{i}}H_{i})^{\prime}M_{\tau_{i}}\omega_{i}]=0$.
Finally, we identify $\tau u_{it}$ since it has zero mean conditional on an appropriate information set $I_{t}=\{\tilde{x}_{it-jk},z_{it-j},L_{i}\leq j\leq U_{i}\}$ with $L_{i}\geq t-T_{i}$ and $U_{i}\leq t_{i}-1$ as a consequence of Assumption 3. This identification strategy was first proposed by Avila Márquez and Kirshnakumar (2023), and it was presented at the 28th International Panel Data Conference at the University of Amsterdam in 2023.
Example.
First-difference transformation: We could follow Baltagi et al. (2002) by first-differencing the structural and reduced-form equations such that we obtain:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\Delta\varepsilon_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (18)
$\Delta x_{1it}=\Delta g_{o}(\tilde{x}_{it},z_{it})+\Delta u_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (19)
Now, we can use the proposed control function approach and augment the structural equation by controlling for the unobservable $\Delta u_{it}$ as follows:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (20)
Next, we set up the following moment conditions:
$\mathbb{E}[(D_{i}H_{i})^{\prime}V_{i}^{-1}D_{i}\omega_{i}]=\textbf{0},$ (21)
where $D_{i}\omega_{i}=D_{i}y_{i}-D_{i}x_{1i}\beta_{1_{o}}-D_{i}\tilde{x}_{i}^{\prime}\beta_{2_{o}}-\rho_{o}D_{i}u_{i}$, $\textbf{0}=[0\quad\cdots\quad 0\quad 0]^{\prime}$, $D_{i}H_{i}=[D_{i}x_{1i}\quad D_{i}\tilde{x}_{i}^{\prime}\quad D_{i}u_{i}]^{\prime}$, and $V_{i}=\sigma^{2}_{\omega}D_{i}D_{i}^{\prime}$.
Finally, we identify $\Delta u_{it}$ since it has zero mean conditional on $\tilde{x}_{it}$, $\tilde{x}_{it-1}$, $z_{it}$, $z_{it-1}$ as a consequence of Assumption 3.
Proposition 1.
Under assumption 3, $\Delta g({\tilde{x}_{it},z_{it}})$ is identified by $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$.
Proof 1.
$\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\mathbb{E}[\Delta g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]+\mathbb{E}[\Delta u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$ (22)
By the linearity property of the expectation operator, we have that the first term is equal to:
$\begin{split}\mathbb{E}[g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-\mathbb{E}[g(\tilde{x}_{it-1},z_{it-1})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\\ g(\tilde{x}_{it},z_{it})-g(\tilde{x}_{it-1},z_{it-1})\end{split}$ (23)
Similarly for the second term, by the Law of Iterated Expectations and Assumption 3:
$\mathbb{E}[u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-\mathbb{E}[u_{it-1}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=0$ (24)
Remark 4.
A similar analysis and conclusion is obtained using a Within transformation which is equivalent to time demeaning the data. In this case, the information set $I_{t}=\{\tilde{x}_{it-j},z_{it-j},-T_{i}+t\leq j\leq t-1\}$.
In Section 6, we propose an estimation method based on the Control Function approach.
6 Super Learner Control Function estimation
As described in the previous section, we can write the transformed structural equation for observation $i$ at period $t$ as follows:
$\tau y_{it}=\tau x_{1it}\beta_{1_{o}}+\tau\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\tau u_{it}+\tau\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (25)
In this transformed equation, the parameters of interest are identified because $\tau x_{1it}$, $\tau\tilde{x}_{it}$ and $\tau u_{it}$ are uncorrelated with error term $\tau\omega_{it}$ under assumption 4.
In addition, it is clear that if $\tau u_{it}$ is observed, we could estimate the three parameters $\beta_{1_{o}}$, $\beta_{2_{o}}$, $\rho_{o}$ using the sample counterparts of the population moment conditions 17. But since $\tau u_{it}$ is not observed, we propose the following estimation procedure:
First Step:
We partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and denote $n_{T,b}=\sum_{i\in S_{b}}T_{i}$. This sample splitting is equal to the one proposed by Emmenegger and Bühlmann (2023) and it was developed without knowledge of it. Alternatively, if the cluster structure in the data is known (individuals exhibit correlation within clusters but not across them, and both cluster membership and the number of clusters are identified) the data should be partitioned accordingly. Then, we estimate $\tau g_{o}(\tilde{x}_{it},z_{it})=\mathbb{E}[\tau x_{1it}|I_{t}]$ using a super learner with partition $S_{b}^{c}=\{i\in S_{g},g\neq b\}$, and we call the estimation $\widehat{\tau g_{o}}^{S_{b}^{c}}$. Then, we obtain the residuals $\widehat{\tau u}_{it}^{S_{b}}=\tau x_{1it}-\widehat{\tau g_{o}^{S_{b}^{c}}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}=\{i\in S_{b}\}$.
Finally, the estimator of $\theta_{o}$ for partition $S_{b}$ is the solution of the sample moment conditions such that:
$\frac{1}{n_{T,b}}\sum_{i\in S_{b}}\phi(O_{i};\hat{\theta}_{b},\widehat{\tau g_{o}}^{S_{b}^{c}})=0,$ (26)
where $\phi(O_{i};\hat{\theta}_{b},\widehat{\tau g_{o}}^{S_{b}^{c}})=(\tilde{H}_{i}^{S_{b}})^{\prime}V_{i}^{-1}M_{\tau_{i}}(y_{i}-\tilde{H_{i}}^{S_{b}}\hat{\theta}_{b})$, $\tilde{H}_{i}^{S_{b}}=[M_{\tau_{i}}x_{1i}\quad M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u}_{i}^{S_{b}}]$, $\widehat{\tau u}_{i}^{S_{b}}$ is the vector stacking the residuals obtained in the first stage for partition $S_{b}$, $\tilde{X}_{i}$ a $T_{i}\times(K-1)$ vector collecting all time observations of $\tilde{X}_{it}$.
This is equivalent to estimate $\theta_{o}$ by performing generalized least squares regression of $\tau y_{it}$ on $\tau x_{1it}$, $\tau\tilde{x}_{it}$ and $\widehat{\tau u}_{it}^{S_{b}}$ or minimizing the following quadratic loss function:
$\hat{\theta}_{b}=\begin{subarray}{c}argmin\\ \theta\end{subarray}Q_{b}(O_{i};\theta,\widehat{\tau g_{o}}^{S_{b}^{c}}),$
with $Q_{b}(O_{i};\theta,\widehat{\tau g_{o}}^{S_{b}^{c}})=\frac{1}{n_{T,b}}\sum_{i\in S_{b}}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}x_{1i}\beta_{1}-M_{\tau_{i}}\tilde{X}_{i}\beta_{2}-\widehat{\tau u_{i}}^{S_{b}}\rho)^{\prime}V_{i}^{-1}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}x_{1i}\beta_{1}-M_{\tau_{i}}\tilde{X}_{i}\beta_{2}-\widehat{\tau u_{i}}^{S_{b}}\rho)$. Importantly, we can just perform ordinary least squares estimation and use a sandwich type variance-covariance matrix estimator for robust inference. In addition, OLS is the most efficient if $M_{\tau_{i}}=D_{i}$, $T=2$, and $\omega_{it}$ is homoskedastic and uncorrelated across both $i$ and $t$.
Second Step:
We average $\hat{\theta}_{b}$ to obtain the final estimator of $\theta_{o}$ as:
$\hat{\theta}_{o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\theta}_{b}.$
Thus, the estimator of our parameter of interest $\beta_{1o}$ is:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\beta}_{b}.$
In order to deal with the dependency of the estimator $\hat{\beta}_{1o}$ on the particular sample split. In order to deal with this issue, we follow Chernozhukov et al. (2018) by proposing to repeate the procedure $ss$ times. The estimates are aggregated either by using the mean or the median. In addition, a correction term is added to the variance estimator of $\hat{\beta}_{1o}$ (See Subsection 8.1.1).
Algorithm 1 Super Learner Control Function estimation
Input: Panel data sets $\{y_{i},X_{i},z_{i}\}_{i\in\{1,2,...,N\}}$ from model satisfying assumptions 1 to 4, a natural number B.
Output: An estimator of the parameter of interest $\beta_{1o}$.
for $ss\in\{1,\dots,SS\}$ do
  Split the individual index set $\{1,2,...,N\}$ into B sets $S_{1},S_{2},...,S_{B}$.
  for $b\in\{1,2,...,B\}$ do
   Estimate the conditional expectation $\mathbb{E}[\tau x_{1it}|I_{t}]$ using a super learner and data corresponding to $S_{b}^{c}$.
   Predict $\mathbb{E}[\tau x_{1it}|I_{t}]$ for $i\in S_{b}$.
   Estimate the residuals $\tau u_{it}$ for $i\in S_{b}$ using the prediction of $\mathbb{E}[\tau x_{1it}|I_{t}]$ of previous step.
   Estimate $\theta_{o}=[\beta_{1o},\beta_{2o},\rho_{o}]^{\prime}$ solving the sample moment conditions 26 for $i\in S_{b}$.
  end for
  Compute $\hat{\theta}_{ss}=\frac{\sum_{1}^{B}\widehat{\theta_{b,ss}}}{B}.$
end for
 Compute $\hat{\theta}_{o}=\frac{\sum_{1}^{SS}\widehat{\theta_{ss}}}{SS}.$
Example.
First-difference transformation (continued):
As described in the previous section, we can write the first-differenced structural equation for observation $i$ at period $t$ as follows:
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (27)
First, we partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and we estimate $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner with partition $S_{b}^{c}$. Then, we get the residuals $\widehat{\Delta u}_{it}^{S_{b}}=\Delta x_{1it}-\widehat{\Delta g_{o}^{S_{b}^{c}}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}$. We obtain the estimator of $\theta_{o}$ for partition $S_{b}$ as the solution of the sample moment conditions such that:
$\frac{1}{n_{T,b}}\sum_{i\in S_{b}}\phi(O_{i};\hat{\theta}_{b},\widehat{\Delta g}_{o}^{S_{b}^{c}})=0,$ (28)
where $\phi(O_{i};\hat{\theta}_{b},\widehat{\Delta g}_{o}^{S_{b}^{c}})=(\tilde{H}_{i}^{S_{b}})^{\prime}V_{i}^{-1}(D_{i}y_{i}-\tilde{H}_{i}^{S_{b}}\hat{\theta}_{b})$, $\tilde{H}_{i}^{S_{b}}=[D_{i}x_{1i}\quad D_{i}\tilde{x}_{i}\quad\widehat{\Delta u}_{i}^{S_{b}}]$, and $\widehat{\Delta u}_{i}^{S_{b}}$ are the residuals obtained in the first stage for partition $S_{b}$.
Finally, we estimate our parameter of interest $\beta_{1o}$ using:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\hat{\beta}_{1o,b}.$
Algorithm 2 Super Learner Control Function Estimation: First-difference transformation
  Input: Panel data sets $\{y_{i},X_{i},z_{i}\}_{i\in\{1,2,...,N\}}$ from model satisfying assumptions 1 to 4, a natural number B.
  Output: An estimator of the parameter of interest $\beta_{1o}$.
  for $ss\in\{1,2,...,SS\}$ do
  Split the individual index set $\{1,2,...,N\}$ into B sets $S_{1},S_{2},...,S_{B}$.
  for $b\in\{1,2,...,B\}$ do
   Estimate the conditional expectation $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner and data correspondig to $S_{b}^{c}$.
   Predict $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ for $i\in S_{b}$.
   Obtain the residuals $\Delta u_{it}$ for $i\in S_{b}$ using the prediction of $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ of previous step.
   Estimate $\theta_{o}=[\beta_{1o},\beta_{2o},\rho_{o}]$ solving the sample moment conditions 26 for $i\in S_{b}$.
  end for
  Compute $\hat{\beta}_{1o,ss}=\frac{\sum_{1}^{B}\widehat{\beta_{1o,b,ss}}}{B}.$
  end for
  Compute $\hat{\beta}_{1o}=\frac{\sum_{1}^{SS}\widehat{\beta_{1o,ss}}}{SS}.$
7 The SLCF control function estimator is a debiased Machine Learning (DML) estimator
In this section, we show that the SLCF estimator is a debiased Machine Learning (DML) estimator. The reason is that the moment condition 17 has a Neyman orthogonal equivalent. Without loss of generality, we assume that there are not covariates such that the model 13 simplifies to:
$\tau y_{it}=\tau x_{1it}\beta_{1_{o}}+\rho_{o}\tau u_{it}+\tau\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (29)
If we stack up the the observations in matrix form, we obtain:
$M_{\tau_{i}}y_{i}=M_{\tau_{i}}x_{1i}\beta_{1_{o}}+\rho_{o}M_{\tau_{i}}u_{i}+M_{\tau_{i}}\omega_{i},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (30)
Now, we premultiply 31 by the projection matrix in the orthogonal space of $\tau u_{i}$ ($A_{i}=I_{i}-M_{\tau_{i}}u_{i}(u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i})^{-1}u_{i}^{\prime}M_{\tau_{i}}^{\prime}$), we obtain:
$A_{i}M_{\tau_{i}}\tau y_{i}=A_{i}M_{\tau_{i}}x_{1i}\beta_{1_{o}}+A_{i}M_{\tau_{i}}\tau\omega_{i},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (31)
Using the transformed model 31, we know that the score for identification of $\beta_{1}$ is:
$\psi(O_{i};\beta_{1},\tau g)=x_{1i}^{\prime}M_{\tau_{i}}^{\prime}A_{i}V_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}.$ (32)
Proposition 2.
The score 32 is Neyman Orthogonal (see appendix for the proof 13.4).
8 Large sample properties of the SLFC estimator
Theorem 1.
Consistency of the SLCF estimator
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, i) $\Omega_{i}$ is positive semi-definite matrix and $\Omega_{i}\mathbb{E}[\phi(O_{i};\theta,\tau g_{o})]=0$ only if $\theta=\theta_{o}$, ii) $\theta\in\Theta$ with $\Theta$ compact, iii) $\phi(O_{i};\theta,\tau g_{o})$ is continuous at each $\theta\in\Theta$, iv) the function class $\mathcal{F}=\{\phi(O_{i};\theta,\tau g_{o}\}$ satisfies a uniform integrability condition $\mathbb{E}[\operatorname*{sup_{\theta\in\Theta}}\left\|\phi(O_{i};\theta,\tau g_{o})\right\|]<\infty$, then $\hat{\theta}_{o}\overset{p}{\to}\theta_{o}$.
The proof is presented in the appendix (Subsection 13.5).
Theorem 2.
Asymptotic normality of the SLCF estimator of $\beta_{1o}$
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, then,
$\sqrt{N_{T}}(\hat{\beta}_{1o}-\beta_{1o})\xrightarrow{d}\mathcal{N}(0,\sigma_{o}^{2}),$ (33)
with $\sigma_{o}^{2}=\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]^{-1}\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}V_{i}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]^{-1}$.
The proof is presented in the appendix (Subsection 13.6).
Theorem 3.
Asymptotic normality of the SLCF estimator of $\theta_{o}$
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, then,
$\sqrt{N_{T}}(\hat{\theta}_{1o}-\theta_{1o})\xrightarrow{d}\mathcal{N}(0,\Sigma_{o}),$ (34)
with $\Sigma_{o}=\mathbb{E}_{P}[H_{i}^{\prime}M_{\tau_{i}}\tilde{V}_{i}^{-1}M_{\tau_{i}}H_{i}]^{-1}\mathbb{E}_{P}[H_{i}^{\prime}M_{\tau_{i}}\tilde{V}_{i}^{-1}\omega_{i}\omega_{i}^{\prime}\tilde{V}_{i}^{-1}M_{\tau_{i}}H_{i}]\mathbb{E}_{P}[H_{i}^{\prime}M_{\tau_{i}}\tilde{V}_{i}^{-1}M_{\tau_{i}}H_{i}]^{-1}$.
The proof is not presented as it follows a similar argument as the proof of Theorem 2.
8.1 Estimator of the variance-covariance matrix
The estimator of the variance-covariance matrix of the SLCF estimator of $\theta_{o}$ is given by:
$\hat{\Sigma}_{o}=\hat{J}_{o}^{-1}\frac{1}{B}\sum_{b=1}^{B}\mathbb{E}_{n,b}[\hat{H^{\prime}}_{i}^{S_{b}}\tilde{V}^{-1}_{i}M_{\tau_{i}}(M_{\tau_{i}}y_{i}^{S_{b}}-\hat{H}_{i}^{S_{b}}\hat{\theta}_{o})(M_{\tau_{i}}y_{i}^{S_{b}}-M_{\tau_{i}}\hat{H}_{i}^{S_{b}}\hat{\theta}_{o})^{\prime}\tilde{V}^{-1}_{i}M_{\tau_{i}}\hat{H}_{i}^{S_{b}}]\hat{J}_{o}^{-1},$ (35)
where:
$\hat{J}_{o}^{-1}=\frac{1}{B}\sum_{b=1}^{B}\mathbb{E}_{n,b}[\hat{H^{\prime}}_{i}^{S_{b}}\tilde{V}^{-1}_{i}\hat{H}_{i}^{S_{b}}].$ (36)
8.1.1 Including sample splitting uncertainty
In order to take into account the sample splitting uncertainty, we can repeate the sample splitting process several times as explained in Section 7 and algorithm 1. We also need to add correction term to the variance-covariance estimator $\hat{\Sigma}_{o}$ given by:
$C=\frac{1}{SS}\sum_{ss=1}^{SS}(\hat{\theta}_{ss}-\hat{\theta}_{o})(\hat{\theta}_{ss}-\hat{\theta}_{o})^{\prime},$ (37)
with $\hat{\theta}_{o}=\frac{1}{SS}\sum_{ss=1}^{SS}\hat{\theta}_{ss}$.
Algorithm 3 Estimator of the variance-covariance of the Super Learner Control Function estimator
Input: Panel data sets $\{y_{i},X_{i},z_{i}\}_{i\in\{1,2,...,N\}}$ from model satisfying assumptions 1 to 4, a natural number B. The selected index set $\{1,2,...,N\}$ for each set $S_{1},S_{2},...,S_{B}$ across all samples splits ($S$) obtained in algorithm 1. The estimated $\hat{\theta}_{ss}$ obtained in algorithm 1.
Output: An estimator of the variance-covariance matrix of $\theta_{1o}$.
for $ss\in\{1,2,...,SS\}$ do
  for $b\in\{1,2,...,B\}$ do
   Estimate $\hat{J}_{o,b}^{-1}=\mathbb{E}_{n,b}[\hat{H^{\prime}}_{i}^{S_{b}}\hat{H}_{i}^{S_{b}}]$.
   Estimate $\mathbb{E}_{n,b}[\hat{H^{\prime}}_{i}^{S_{b}}M_{\tau_{i}}(M_{\tau_{i}}y_{i}^{S_{b}}-\hat{H}_{i}^{S_{b}}\hat{\theta}_{ss})(M_{\tau_{i}}y_{i}^{S_{b}}-M_{\tau_{i}}\hat{H}_{i}^{S_{b}}\hat{\theta}_{ss})^{\prime}M_{\tau_{i}}\hat{H}_{i}^{S_{b}}]$.
  end for
  Compute $\hat{\Sigma}_{s}.$
end for
 Compute $C=\frac{1}{SS}\sum_{ss=1}^{SS}(\hat{\theta}_{ss}-\hat{\theta}_{o})(\hat{\theta}_{ss}-\hat{\theta}_{o})^{\prime}.$
 Compute $\hat{\Sigma}_{o}=\frac{1}{SS}\sum_{ss=1}^{SS}(\hat{\Sigma}_{s})+C.$
9 Naive plugin IV estimation is inconsistent and it is not numerically equivalent to SLCFE
The Super Learner Control function estimator is not numerical equivalent to an IV estimator that plugs in the estimated transformed nuisance parameter $\tau g_{o}(\tilde{x}_{it},z_{it})$. Without loss of generality, we assume that exogenous covariates are not present (model 1 to 3). As described in Section 5, one could be tempted to estimate the parameter of insterest by using the naive IV estimator:
$\hat{\beta}_{IV,1_{o}}=\sum_{b}^{B}\frac{1}{B}(\sum_{i\in S_{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i})^{-1}(\sum_{i\in S_{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}).$ (38)
Theorem 4.
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, Naive plugin IV estimation is not numerically equivalent to SLCFE when $g_{o}(.)$ is nonlinear
Proof 2.
Assuming that $B=1$, we can write the estimators as:
$\hat{\beta}_{IV,1_{o}}=\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big),$ (39)
$\hat{\beta}_{SLCFE,1_{o}}=\Big(\sum_{i}(M_{\tau_{i}}x_{1i})^{\prime}\hat{A}_{i}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i}(M_{\tau_{i}}x_{1i})^{\prime}\hat{A}_{i}M_{\tau_{i}}y_{i}\Big),$ (40)
where $\hat{A}_{i}=I_{i}-\widehat{\tau u_{i}}(\widehat{\tau u_{i}}^{\prime}\widehat{\tau u_{i}})^{-1}\widehat{\tau u_{i}}^{\prime}$.
Because $\hat{A}_{i}M_{\tau_{i}}x_{1i}\neq\widehat{\tau g_{o}}(z_{i})$ , the two estimators are not equivalent. They are numerically equivalent if and only if $\tau g_{o}(z)\in\text{col}(\tau z)$.
Theorem 5.
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold in particular that $\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}$ with $\delta_{N}\sqrt{N_{T}}\to\infty$, $\{\delta_{N}\}$ a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$ , the Naive plugin IV estimator is inconsistent.
Proof 3.
The naive plugin IV estimator with sample splitting is given by:
$\hat{\beta}_{IV,1_{o}}=\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big).$ (41)
Replacing $M_{\tau_{i}}y_{i}$ by $\tau g_{o}(z_{i})\beta_{1_{o}}+M_{\tau_{i}}u_{i}\beta_{1_{o}}+M_{\tau_{i}}\varepsilon_{i}$, adding and substracting $\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})\beta_{1_{o}}$, it can be written as:
$\hat{\beta}_{1,o}=\beta_{1,o}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}(\xi_{i}^{S_{b}^{c}}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (42)
with $\xi_{i}^{S_{b}^{c}}=\tau g_{o}(z_{i})-\widehat{\tau g_{o}}(z_{i})^{S_{b}^{c}}$.
Now, adding and substracting $\tau g_{o}(z_{i})$ to $\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})$ we obtain that:
$\hat{\beta}_{1,o}=\beta_{1,o}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}(\xi^{S_{b}^{c}}_{i}+\tau g_{o}(z_{i}))^{\prime}(\xi^{S_{b}^{c}}_{i}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (43)
The second term does not converge to zero at rate faster than $\sqrt{N_{T}}$ because $\tau g_{o}(z_{i})$ is not centered at 0. Thus, the naive plugin estimator presents a first order bias term that does not vanish using sample-splitting. Thus, the naive plugin IV estimator is inconsistent.
10 Monte Carlo simulation
We test the performance of the proposed estimation methods using a Monte Carlo simulation experiment. As competing methods, we consider Within OLS, and Within 2SLS estimators including polynomial transformations of the covariates and instruments of degree 1 and 5.
10.1 Settings
10.1.1 Linear structural equation and non-linear secondary equation
We generate 100 samples from a triangular simultaneous model with a linear structural equation and a non-linear reduced-form equation with $N=1000$ and $T=2$. Both equations present additive disturbance terms, and individual-specific effects. The structural equation presents one endogenous regressor $x_{1it}$ and one exogenous covariate $x_{2it}$.
$\begin{split}y_{it}=\beta_{1}x_{1it}+\beta_{2}x_{2it}+\alpha_{i}+\epsilon_{it},\end{split}$ (44)
where:
$x_{2it}=\alpha_{i}+\zeta_{it},\quad with\quad\zeta_{it}\sim U(-2,2),$ (45)
$z_{it}=\alpha_{i}+\nu_{it},\quad with\quad\nu_{it}\sim U(-2,2),$ (46)
$x_{1it}=g(x_{2it},z_{it})+\alpha_{i}+u_{it},$ (47)
$u_{it}\sim U(-1,1),$ (48)
$\epsilon_{it}=0.9*u_{it}+\tilde{\zeta_{it}},\quad with\quad\tilde{\zeta}_{it}\sim U(-1,1).$ (49)
$\alpha_{i}\sim U(-1,1).$ (50)
The unknown function $g(x_{2it},z_{it})$ is similar to the ones proposed by Guo et al. (2022), and depends on the parameter $a$ that controls the nonlinearity in the functional form and the strength of the linear correlation between $x_{1it}$ and $z_{it}$. As in Guo et al. (2022), a larger $a$ causes higher nonlinearity and lower linear correlation between $x_{1it}$ and $z_{it}$. However, we cannot compare our results to theirs because they consider a model for cross-sectional data.
Table 1: Monte Carlo Experiment
Scenario $g(\cdot)$ $\beta_{1}$ $dim(x_{2})$ $\beta_{2}$
1 $a\cdot|z_{it}|-2tanh(x_{2it})+1/a*z_{it}$ 1 1 1
((a))
((b))
Figure 1: a) Relationship between $\Delta x_{1it}$ and $z_{it}$ with $a=10$, b) Relationship between $\Delta x_{1it}$ and $z_{it}$ with $a=1$
10.2 The results
In this section, we present the results of the simulation experiment. In figure 2, we present the average estimated parameter of interest $\beta_{1o}$ for different values of $a$. The methods are Within OLS (WOLS), Within 2SLS (W2SLS) using polynomial of degree 1 of the covariate and the instrumental variables, Within 2SLS using polynomial of degree 5 of the covariate and the instrumental variables (W2SLS_polynomial), the Super Learner Control Function estimator on the First-Differenced data (FDSLCF), and the Super Learner Control Function estimator on the Within-transformed data (WSLCF). We also plot the p-value of the estimated control functions using the first-differenced and the within transformed data. As $a$ is larger, the nonlinearity in the reduced form equation is more important. This translates in a lower correlation between the endogenous regressor and the instrumental variables. As a result, the Within-2SLS estimators present a decay in their performance. In contrast, the Super Learner Control Function estimators remain stable around the true value of the parameter of interest $\beta_{1o}=1$.
In panel (a) of figure 3, we present the boxplot of the estimated parameter of interest $\beta_{1o}$ for a highly nonlinear reduced form equation (DGP 1 with $a=5$), and in panel (b) for a linear reduced form equation (DGP 1 with $a=1$). The estimated parameters with the Super Learner Control Function estimators are tightly concentrated around the true value of the parameter of interest. In contrast, the Within estimates present wide variation around the true value.
For the estimation of the nuisance parameters, we used a Super Learner. The base learners used for the Super Learner are a linear model, a neural network, and a random forest. The architecture of the neural network is one hidden layer with two neurons, the activation function is the sigmoid one, and the output function is logistic, the regularization is 0, and the maximum number of iterations is 100. The random forest is composed of 100 trees with minimum leaf size of 5, and two variables to select at each split.
Figure 2: Average estimates of $\beta_{1o}$ for 100 samples simulated with different values of $a$, $N=1000$, $T=2$
((a))
((b))
Figure 3: a) Estimates of $\beta_{1o}$ for a DGP highly nonlinear (a=5), b) Estimates of $\beta_{1o}$ for a DGP highly linear (a=1), $N=1000$, $T=2$
Table 2: Coverage results: Nominal coverage 95%
N T a $S$ K $\beta_{1o}$ WOLS W2SLS W2SLS polynomial FDCF WCF
1000 2 1 10 5 1 100 100 100 71 76
1000 2 2 10 5 1 100 100 100 94 90
1000 2 3 10 5 1 100 100 100 99 99
1000 2 4 10 5 1 100 100 100 98 100
1000 2 5 10 5 1 100 100 100 97 100
1000 2 6 10 5 1 100 100 100 97 100
1000 2 7 10 5 1 100 100 100 95 100
1000 2 8 10 5 1 100 100 100 97 100
1000 2 9 10 5 1 100 100 100 96 100
1000 2 10 10 5 1 100 100 100 95 100
1000 2 1 100 5 1 100 100 100 79 81
1000 2 2 100 5 1 100 100 100 95 96
1000 2 3 100 5 1 100 100 100 95 96
1000 2 4 100 5 1 100 100 100 96 97
1000 2 5 100 5 1 100 100 100 99 100
1000 2 6 100 5 1 100 100 100 96 100
1000 2 7 100 5 1 100 100 100 97 100
1000 2 8 100 5 1 100 100 100 99 100
1000 2 9 100 5 1 100 100 100 100 100
1000 2 10 100 5 1 100 100 100 99 100
11 Empirical Application: Air Pollution and Educational Outcomes in the U.S. (Revisited)
As an illustration, we estimate the causal effect of air pollution on student performance in the United States. The data on air pollution are obtained from Deryugina et al. (2019). This dataset presents daily data on air pollutants: PM 2.5, ozone ($O_{3}$), carbon monoxide (CO), sulfur dioxide ($SO_{2}$), nitrogen dioxide ($NO_{2}$). It also contains data on wind direction, wind speed, and weather conditions at county level. We aggregate the data to yearly level by averaging across days. The data on student performance are obtained from Stanford Educational Data Archive (SEDA, Fahle et al. 2024) for county-yearly level. We merge both datasets by county-year indicator. The final dataset is a panel data set with 787 counties for the period 2009 to 2013. We transform the data using the first-difference and within transformations.
We specify the following linear regression for county $d$, subject $s$, cohort $c$ at year $t$:
$y_{scdt}=\beta PM2.5_{dt}+X_{dt}^{\prime}\gamma+\psi_{s}+\phi_{c}+\alpha_{d}+\tilde{\alpha}_{g}+\epsilon_{dst},$ (51)
where $y_{dst}$ represents the average grade of subject $s$, cohort $c$, county $d$, subject $s$ at year $t$. $PM2.5_{dt}$ represents the yearly average of $PM2.5$ concentration at county $d$ in year $t$, $X_{dt}$ is a vector of control variables at county-year level including average school characteristics, social characteristics, and weather variables, $\phi_{c}$ corresponds to cohort specific effects, $\tilde{\alpha}_{g}$ grade specific effects, and $\psi_{s}$ corresponds to subject specific effects. $\alpha_{d}$ represents county fixed effects.
As mentioned before, $PM2.5_{dt}$ is endogenous. Thus, following Deryugina et al. (2019) we use wind direction as an instrument of air pollution along with a nonlinear first-stage equation as follows:
$PM25_{dt}=h(WD_{dt},X_{dt},\psi_{s},\phi_{c},\tilde{\alpha}_{g})+\eta_{d}+\varepsilon_{dt},$ (52)
with $WD_{dt}$ is the yearly average wind direction in county $d$ in year $t$, $\eta_{d}$ are county specific effects.
Using a First-difference transformation, we estimate that an increase in the concentration of PM 2.5 causes a reduction in test scores of 0.976 standard deviations that is statistically significant at a 5% significance level. The base learners used for the Super Learner are the mean, a linear model, a neural network, and a random forest. The architecture of the neural network is one hidden layer with two neurons, the activation function is the sigmoid one, and the output function is linear, the regularization is 0, and the maximum number of iterations is 100. The random forest is composed of 100 trees with minimum leaf size of 5, and two variables to select at each split. The number of sample splits is equal to 10, and the number of folds is 5. We use $\tilde{V}_{i}=I_{i}$ which is the most efficient in the case of the Within transformation. Inference is still valid as we used a sandwich type estimator of the variance-covariance matrix of the coefficients of the structural model.
Table 3: Estimated effect of air pollution (PM 2.5 concentration) on student performance
WOLS W2SLS FD SLCF W SLCF
-0.072 -0.844 -0.979 -0.859
(0.011) (0.762) (0.101) (0.085)
Note: US counties, (N = 787).
12 Conclusions
This paper proposes a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation to deal with the issue of weak instrumental variables due to nonlinearities in panel data settings. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. We estimate the control function with a super learner using sample splitting. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. The estimator is consistent and asymptotically normal achieving a parametric rate of convergence. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the estimator perform well, provided that we can learn accurately the nuisance parameter in the first stage.
13 Appendix
13.1 Notation
A set $\{1,2,...,A\}$ is denoted as $[A]$. $|\cdot|$ refers to the cardinality of a set. $\left\|\cdot\right\|_{L^{p}}$ is the $L^{p}$ norm defined on the space $L^{p}(\Omega,P)$. $\left\|\cdot\right\|_{p}$ is the p-norm.
13.2 Definitions
We follow Emmenegger and Bühlmann (2023) and provide the following definitions. Each individual $i$ presents $T_{i}$ observations. We assume that $T_{i}$ is uniformly bounded such that $T_{i}<T_{max}$. The total number of observations is denoted by $N_{T}=\sum_{i}T_{i}$. The number of sample splits $B$ is not random and independent of $N$. The data is defined as $\{O_{i}=(X_{i},z_{i},y_{i})\}_{i\in[N]}$ with $X_{i}=[x_{1i},\tilde{X}_{i}]$, and denote the probability distribution $P$ of the grouped data $O_{i}$. $\{\mathcal{P}_{N}\}_{N\geq 1}$ is a sequence of probability distributions. A partition of the set $[N]$ is $S_{1}$, $S_{2}$, …, $S_{B}$. The total number of observations belonging to a partition $S_{b}\in[N]$ is equal to $n_{T,b}=\sum_{i\in S_{b}}T_{i}$. We assume that the subsets $S_{b}$ have similar size such that $Bn_{T,b}=N_{T}+o(1),\quad\forall b\in[B]$. We denote $S_{b}^{c}=\{O_{i}=(y_{i},X_{i},z_{i})\}_{i\in S_{b}^{c}}$, and the estimator of the nuisance parameter $\widehat{\tau g}_{o}^{S_{b}^{c}}$ with data from $S_{b}^{c}$. The empirical mean $\mathbb{E}_{n_{T,b}}=\frac{1}{n_{T_{b}}}\sum_{i\in S_{b}}\phi(O_{i};\theta,\tau g)$. We define $\{\delta_{N}\}_{N\geq B}$ as a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$, and $\{\delta_{N}\}\geq N^{-1/4}$.
13.2.1 Population Loss Function
We define the population following population loss function:
$Q(\theta,\tau g_{o})=\mathbb{E}[(M_{\tau_{i}}(y_{i}-H_{i}\theta))^{\prime}V_{i}^{-1}M_{\tau_{i}}(y_{i}-H_{i}\theta)],$
where $V_{i}=\mathbb{E}[(M_{\tau_{i}}\omega_{i})(M_{\tau_{i}}\omega_{i})^{\prime}|H_{i}]$, and $\theta=[\beta_{1},\quad\beta_{2}^{\prime},\quad\rho]^{\prime}$. $V_{i}$ is a scalar if $T_{i}=2,\forall i\in[N]$, $M_{\tau_{i}}=D_{i}$, and the error term $\omega_{it}$ is homoskedatisc and uncorrelated across $i$ and $t$.
13.2.2 Empirical Loss Function
We define the population following loss function:
$L_{N}(\theta,\tau g_{o})=\sum_{i}^{N}\frac{1}{N}[(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)].$
13.2.3 Population moment condition
We define the population moment condition:
$m(O_{i};\theta,\tau g_{o})=\mathbb{E}[(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}M_{\tau_{i}}H_{i}]$
with $m(O_{i};\theta_{o},\tau g_{o})=0$.
13.2.4 Score function
The score function is $\phi(O_{i};\theta,\tau g_{o})=(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}M_{\tau_{i}}H_{i}$.
13.2.5 Sample moment condition
We define the sample moment condition:
$\hat{m}(O_{i};\theta,\tau g_{o})=\frac{1}{N}\sum_{i}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}M_{\tau_{i}}H_{i}.$
13.3 Assumptions
Assumption 5.
$\theta\in\Theta$, the set $\Theta$ is bounded, contains $\theta_{o}$.
This assumption is similar to assumption 4.1 of Emmenegger and Bühlmann (2023).
Assumption 6.
The set $\mathcal{G}$ consists P-integrable functions $\tau g$ with pth moment that exists and it contains $\tau g_{o}$, and
$\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}.$
Assumption 7.
For all N, all $i\in[N]$, all $P\in\mathcal{P}_{N}$, all $b\in[B]$, and $p>8$ we have:
  1. 7.1
    At the true $\theta_{0}$ and the true $g_{0}$, the data $\{O_{i}=(W_{i},X_{i},Z_{i},Y_{i})\}_{i\in[N]}$ satisfies the identifiability condition.
    $\mathbb{E}_{P}\Big[\mathbb{E}_{n_{T,b}}[\phi(O_{i};\theta_{o},\tau g_{o})]\Big]=0.$
  2. 7.2
    $\mathbb{E}_{P}\Big[\mathbb{E}_{n_{T,b}}[\phi(O_{i};\theta_{o},\tau g_{o})\phi(O_{i};\theta_{o},\tau g_{o})^{\prime}]\Big]=\Omega_{i}$ with $\Omega_{i}$ a semi positive-definite matrix, with eigenvalues bounded away from 0 and finite, with the minimum eigenvalue $\tilde{\lambda}_{T}>0$, and the maximum eigenvalue $\tilde{\lambda}_{1}<\infty$.
  3. 7.3
    $V_{i}$ is a semipositive-definite matrix, with eigenvalues bounded away from 0 and finite, with the minimum eigenvalue $\lambda_{T}>0$, and the maximum eigenvalue $\lambda_{1}<\infty$.
  4. 7.4
    There exists finite real constants $C_{1}$$C_{2}$, $C_{3}$, $C_{4}$ satisfying $y_{i}||_{P,p}\leq C_{1}$, $z_{i}||_{P,p}\leq C_{2}$, $x_{1,i}||_{P,p}\leq C_{3}$, $\tilde{x}_{i}||_{P,p}\leq C_{4}$ for all $k>2$.
Assumption 7.1 states that the true parameter vector is identifiable. Assumption 7.4 state that the data has finite moments.
This assumption is equivalent to Assumption 5.1 of Emmenegger and Bühlmann (2023), and Newey and McFadden (1994).
13.4 Proof of the Orthogonal Score
Without loss of generality, we assume that $V_{i}=I_{i}$. Then, we define the Gateaux derivative of the score function $\mathbb{E}[\psi(O_{i};\beta_{1o},\tau g_{o})]$ as follows:
$D_{h}[\tau g-\tau g_{o}]:=\lim_{h\to 0}\frac{\mathbb{E}_{P}[\psi(O_{i};\beta_{1o},\tau g_{o}+h(\tau g-\tau g_{o}))]-\mathbb{E}_{P}[\psi(O_{i};\beta_{1o},\tau g_{o})]}{h}.$ (53)
Then, the Gateaux derivative is given by:
$D_{h}[\tau g-\tau g_{o}]=\lim_{h\to 0}\frac{\mathbb{E}_{P}[x_{1i}M_{\tau_{i}}^{\prime}A_{i,h}M_{\tau_{i}}\omega_{i}]-\mathbb{E}_{P}[x_{1i}M_{\tau_{i}}^{\prime}A_{i}M_{\tau_{i}}\omega_{i}]}{h}.$ (54)
with:
$A_{i,h}=I_{i}-\frac{M_{\tau_{i}}(u_{i}-hg(\tilde{x}_{i},z_{i}))(u_{i}-hg(\tilde{x}_{i},z_{i}))^{\prime}M_{\tau_{i}}^{\prime}}{(u_{i}-hg(\tilde{x}_{i},z_{i}))^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}(u_{i}-hg(\tilde{x}_{i},z_{i}))},$ (55)
$A_{i}=I_{i}-\frac{M_{\tau_{i}}u_{i}u_{i}^{\prime}M_{\tau_{i}}^{\prime}}{u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}}.$ (56)
Now, we can re-write $A_{i,h}$ as follows:
$A_{i,h}=I_{i}-\frac{M_{\tau_{i}}(u_{i}u_{i}^{\prime}-hg(\tilde{x}_{i},z_{i})u_{i}^{\prime}-u_{i}hg(\tilde{x}_{i},z_{i})^{\prime}+h^{2}g(\tilde{x}_{i},z_{i})g(\tilde{x}_{i},z_{i})^{\prime})M_{\tau_{i}}^{\prime}}{u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}-u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}hg(\tilde{x}_{i},z_{i})-hg(\tilde{x}_{i},z_{i})^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}+h^{2}g(\tilde{x}_{i},z_{i})^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}g(\tilde{x}_{i},z_{i})},$ (57)
such that recalling terms, we obtain:
$A_{i,h}=I_{i}-\frac{M_{\tau_{i}}(u_{i}u_{i}^{\prime}-B(h))M_{\tau_{i}}^{\prime}}{u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}-a(h)},$ (58)
Substracting $A_{i}$ from $A_{i,h}$, we obtain:
$A_{i,h}-A_{i}=\frac{M_{\tau_{i}}(B(h))M_{\tau_{i}}^{\prime}u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}-a(h)M_{\tau_{i}}u_{i}u_{i}^{\prime}M_{\tau_{i}}^{\prime}}{[u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}-a(h)]u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}}$ (59)
Then, we obtain that the Gateaux derivative evaluted at $h=0$ is equal to 0 by L.I.E., assumptions 3 and 4:
$D_{h}[\tau g-\tau g_{o}]=\mathbb{E}_{P}[x_{1i}M_{\tau_{i}}^{\prime}\frac{M_{\tau_{i}}(-g(\tilde{x}_{i},z_{i})u_{i}^{\prime}-u_{i}g(\tilde{x}_{i},z_{i})^{\prime})M_{\tau_{i}}^{\prime}u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}+2g(\tilde{x}_{i},z_{i})^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i}M_{\tau_{i}}u_{i}u_{i}^{\prime}M_{\tau_{i}}^{\prime}}{(u_{i}^{\prime}M_{\tau_{i}}^{\prime}M_{\tau_{i}}u_{i})^{2}}M_{\tau_{i}}\omega_{i}]=0.$ (60)
13.5 Consistency
The proof of consistency of $\hat{\theta}_{o}$ follows Emmenegger and Bühlmann (2023), and Newey and McFadden (1994).
Proof 4.
Proof of Theorem 1
Condition i) is satisfied by Assumptions 7.1, and 7.2, condition ii) is satisfied by Assumption 5, condition iii) and iv) are satisfied by lemma 12, lemma 15 of Emmenegger and Bühlmann (2023), and lemma 1. This implies that the estimator $\hat{\theta}$ is consistent (Emmenegger and Bühlmann, 2023).
Lemma 1.
We prove that lemma 13 of Emmenegger and Bühlmann (2023) holds in our set up.
Proof 5.
$\begin{split}\phi(O_{i};\theta,\tau g)-\phi(O_{i};\theta,\tau g_{o})&=(M_{\tau_{i}}y_{i}-\tilde{H}_{i}\theta)^{\prime}V_{i}^{-1}\tilde{H}_{i}-(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}M_{\tau_{i}}H_{i}\\ &=(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}[\tilde{H}_{i}-M_{\tau_{i}}H_{i}]+[(\tilde{H}_{i}-M_{\tau_{i}}H_{i})\theta)]^{\prime}V_{i}^{-1}[(\tilde{H}_{i}-M_{\tau_{i}}H_{i})],\end{split}$ (61)
Now, notice that $(\tilde{H}_{i}-M_{\tau_{i}}H_{i})=[M_{\tau_{i}}x_{1i},M_{\tau_{i}}\tilde{X}_{i},\widehat{\tau u}_{i}]-[M_{\tau_{i}}x_{1i},M_{\tau_{i}}\tilde{X}_{i},M_{\tau_{i}}u_{i}]=[0,0,\widehat{\tau u}_{i}-M_{\tau_{i}}u_{i}]=[0,0,\xi_{i}]$. Then, we are left with:
$(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}.$ (62)
With $P\in\{\mathcal{P}_{N}\}_{N\geq 1}$, then:
$\begin{split}\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\mathbb{E}_{n_{T,b}}[(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]]\right\|&\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}]\right\|+\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}M_{\tau_{i}}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\theta\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\xi_{i}^{\prime})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\lambda_{1}\sum_{t}(q_{t}^{\prime}M_{\tau_{i}}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}[\lambda_{1}\left\|\theta\mathbb{E}_{P}\sum_{t}(q_{t}^{\prime}\xi_{i})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\lambda_{1}\left\|M_{\tau_{i}}y_{i}\right\|_{P,2}\left\|\xi_{i}\right\|_{P,2}\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\lambda_{1}\left\|\theta\right\|\left\|\xi_{i}\right\|_{P,2}\end{split}$ (63)
Since $\Theta$ is bounded, $\left\|\tau g-\tau g_{o}\right\|_{P,2}<\delta_{N}$, the last expression is bounded by $\delta_{N}$ by assumptions 5, and 7.
13.6 Asymptotic Normality
Proof 6.
In order to prove the asymptotic normality of the SLCF estimator of $\beta_{1o}$, we write the estimator as:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\Big(argmin_{\beta_{1}\in\mathbb{R}}\frac{1}{n_{T,b}}\sum_{i\in S_{b}}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}x_{1i}\beta_{1}-M_{\tau_{i}}\tilde{X}_{i}^{\prime}\beta_{2}-\widehat{\tau u_{i}}^{S_{b}}\rho)^{\prime}\hat{A}_{i}^{S_{b}}V_{i}^{-1}\hat{A}_{i}^{S_{b}}(M_{\tau_{i}}y_{i}-M_{\tau_{i}}x_{1i}\beta_{1}-M_{\tau_{i}}\tilde{X}_{i}^{\prime}\beta_{2}-\widehat{\tau u_{i}}^{S_{b}}\rho)\Big),$ (64)
with $\hat{A}_{i}^{S_{b}}=I_{i}-[M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]\big([M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]^{\prime}[M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]\big)^{-1}[M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]^{\prime}$.
Such that:
$\hat{\beta}_{1o}=\frac{1}{B}\sum_{b=1}^{B}\Big(\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}y_{i}\Big),$ (65)
where we use $\tilde{V}_{i}$ because $\sigma_{\omega^{2}}$ is cancelled out, recalling that $V_{i}=\mathbb{E}[(M_{\tau_{i}}\omega_{i})(M_{\tau_{i}}\omega_{i})^{\prime}|H_{i},z_{i}]=\sigma^{2}_{\omega}M_{\tau_{i}}M_{\tau_{i}}^{\prime}=\sigma^{2}_{\omega}\tilde{V}_{i}$.
This is equal to:
$\sqrt{N_{T}}(\hat{\beta}_{1o}-\beta_{1o})=\sum_{b=1}^{B}\Big(\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\frac{\sqrt{N_{T}}}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}(M_{\tau_{i}}u_{i}\rho_{o}+M_{\tau_{i}}\omega_{i})\Big).$ (66)
Since $\hat{A}_{i}^{S_{b}}=A_{i}+o_{p}(1)$ by Lemma 2, we have that:
$\begin{split}\frac{\sqrt{N_{T}}}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}(M_{\tau_{i}}u_{i}\rho_{o}+M_{\tau_{i}}\omega_{i})=&\frac{\sqrt{N_{T}}}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}(M_{\tau_{i}}u_{i}\rho_{o}+M_{\tau_{i}}\omega_{i})+o_{p}(1)\\ =&\frac{\sqrt{N_{T}}}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}+o_{p}(1)\end{split}$ (67)
We define:
$\sigma^{2}_{1i}=\mathbb{E}_{P_{N}}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\omega_{i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}],$ (68)
$\bar{\sigma^{2}_{1}}=\frac{1}{N_{T}}\sum_{i}^{N}\sigma^{2}_{1i}.$ (69)
Now, we verify the Lindeberg condition:
$\begin{split}&\frac{1}{\bar{\sigma^{2}_{1}}N_{T}}\sum_{i}^{N}\mathbb{E}_{P_{N}}\Big[\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|^{2}\mathbbm{1}_{\{\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|\geq\epsilon\sqrt{N_{T}\bar{\sigma^{2}_{1}}}\}}\Big]\\ &\leq\frac{1}{\bar{\sigma^{2}_{1}}N_{T}}\sum_{i}^{N}\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|_{P_{N}}^{2}P_{N}\Big(\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|\geq\epsilon\sqrt{N_{T}\bar{\sigma^{2}_{1}}}\Big)\\ &\leq\frac{1}{\bar{\sigma^{2}_{1}}N_{T}}\sum_{i}^{N}\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|_{P_{N}}^{2}\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|_{P_{N}}\frac{1}{\epsilon\sqrt{N_{T}\bar{\sigma^{2}_{1}}}}\\ &\leq\frac{1}{\bar{\sigma^{2}_{1}}N_{T}}\sum_{i}^{N}\|x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\|_{P_{N}}^{3}\frac{1}{\epsilon\sqrt{N_{T}\bar{\sigma^{2}_{1}}}}\\ &\leq\frac{1}{\sqrt{N_{T}}}\rightarrow 0\text{ as }N_{T}\rightarrow\infty\forall\epsilon>0.\end{split}$ (70)
by Markov inequality, and Assumption 7.2. Then, by the Lindeberg-Feller Central Limit Theorem we have that:
$\bar{\sigma^{2}_{1}}^{-1/2}\frac{1}{\sqrt{N_{T}}}\sum_{i}x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}\omega_{i}\xrightarrow{d}\mathcal{N}(0,1).$ (71)
We define:
$\begin{split}\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}=\mathbb{E}_{n_{T},b}[x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}]\end{split}$ (72)
Also, we have that:
$\begin{split}\mathbb{E}_{n_{T},b}\Big[x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}-\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]\Big]\\ =\mathbb{E}_{n_{T},b}\Big[x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}-\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]\Big]+\\ \mathbb{E}_{n_{T},b}\Big[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}-\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}V_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]\Big]=o_{P}(1).\end{split}$ (73)
by assumptions 6, 7.4, and $\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}=\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}+o_{P}(1)$ by Lemma 2. Then,
$\frac{1}{B}\sum_{b}\frac{1}{n_{T,b}}\sum_{i\in S^{b}}x_{1i}^{\prime}M_{\tau_{i}}\hat{A}_{i}^{S_{b}}\tilde{V}_{i}^{-1}\hat{A}_{i}^{S_{b}}M_{\tau_{i}}x_{1i}=\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}].$ (74)
Thus,
$\sqrt{N_{T}}(\hat{\beta}_{1o}-\beta_{1o})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}_{o}),$ (75)
with $\sigma^{2}_{o}=\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]^{-1}\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}V_{i}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]\mathbb{E}_{P}[x_{1i}^{\prime}M_{\tau_{i}}A_{i}\tilde{V}_{i}^{-1}A_{i}M_{\tau_{i}}x_{1i}]^{-1}$.
Lemma 2.
Notice that $M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]=[M_{\tau_{i}}\tilde{X}_{i}\quad M_{\tau_{i}}u_{i}+M_{\tau_{i}}g(\tilde{x}_{i},z_{i})-\widehat{M_{\tau_{i}}g(\tilde{x}_{i},z_{i})}^{S_{b}}]$, then by Assumption 7.4 we have that:
$M_{\tau_{i}}\tilde{X}_{i}\quad\widehat{\tau u_{i}}^{S_{b}}]=[M_{\tau_{i}}\tilde{X}_{i}\quad M_{\tau_{i}}u_{i}]+o_{p}(1).$ (76)
The linear map $a(L_{i})=I-L_{i}(L_{i}^{\prime}L_{i})^{-1}L_{i}^{\prime}$ is a continuous function since matrix addition, multiplication and inversion are continous functions.
Then, by the Continous Mapping Theorem for convergence in probability we have that $\hat{A}_{i}^{S_{b}}=A_{i}+o_{p}(1)$.